### v1

# Matching Non-Identical Objects

Abstract
Not identical but similar objects are everywhere in the world. Examples include four-legged animals such as dogs and cats, cars of different models, akin flowers in various colors, and countless others. In this study, we address a novel task of matching such non-identical objects. We propose a simple weighting scheme of descriptors that enhance various sparse image matching methods, which are originally designed for matching identical objects captured from different perspectives, and achieve semantically robust matching. The experiments show successful matching between non-identical objects in various cases including domain shift. Further, we present a first evaluation of the robustness of the image matching methods under common corruptions, which is a sort of domain shift, and the proposed method improves the matching in this case as well.
1 Introduction
Figure 1: Non-identical object matching is achieved densely and consistently by adding our method between SuperPoint [7] and LightGlue [17]. Object correspondences can be found even when the design is different as in (a), the classes are discrepant as in (b), and there are several similar objects as in (c). Objects are robustly matched to each other even with domain shifts in pictures and photos as in (d) and corruption as in (e) and (f).
We humans are able to find fine-grained visual correspondences between two objects even when they look similar but not identical or when one is sketched and another appears in a photo. Thanks to this ability, for example, we can identify a cat in gloomy backstreet from a photo taken in a bright room or assemble components to build a shelf based on its perspective drawing. In computer vision words, this needs fine-grained matching between feature points of two objects under image corruptions, class shift, and domain shift. The classical image matching task [30, 36, 17, 11] only addresses a special case; objects to be matched are assumed to be identical and captured from different perspectives. Recent Transformer-based dense matching methods show a great performance on such a matching task [36, 26, 48, 14, 8, 46, 4, 45], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [30, 3, 44, 27, 23, 11, 19]. However, they do not perform well once the identicality assumption is not violated and/or objects are captured in harsh environments or different domains (cf. Fig. 2).
In this study, we address a novel task of non-identical object matching, where objects to be matched are not necessarily identical. For example, the objects to be matched can be clocks of different designs or animals of different species. Further, the non-identicality naturally induces the necessity of robustness against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s. drawing) because two objects may be captured in different environments and domains, as shown in Fig. 1. A crucial challenge of achieving non-identical object matching is that we cannot simply resort to a supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects. At least, to our knowledge, there are no such datasets. Interestingly, we discovered a simple plug-and-play weighting module can extend off-the-shelf sparse matching methods to the non-identical object matching. Namely, given a pair of images, sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints. Our weighting module can be inserted before the feature matching model to weights the descriptors using semantic information. Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [32] to obtain weights. Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in non-identical object matching task. The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features.
Our experiments demonstrate that this simple approach significantly boosts the robustness of sparse matching methods, including X and Y, in two tasks: non-identical object matching and robust image matching. In the former task, we evaluate the matching results with the image pairs (a) of the same class, (b) of different classes, (c) with domain shift, and (d) of the same appearance. The sparse matchers alone resulted in scattered mismatching between non-corresponding areas, but with the addition of our method, fine-grained and consistent matching is achieved. In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [36] dataset under common corruptions [13] in terms of the AUC score of standard relative pose estimation. The proposed method increases the AUC scores with various types of common corruptions, and exceeds a dense matching method LoFTR [36] in the average case.
To summarize, our contributions are as follows:
  • We propose non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images even under image corruptions and/or between different domains. This task can be considered a generation of the classical image matching task.
  • We propose a simple plug-and-play weighting module that enhances the low-level features using semantic information and successfully extends various sparse matching methods to non-identical object matching.
  • We present a first evaluation of the robustness in image matching under common corruptions. By adding corruptions to both input images, the robustness of corruptions itself is measured. By adding them to only one of the input images, the robustness against a domain shift is measured. For both cases, the proposed method is shown useful.
2 Related Work
Image matching [30, 36, 17, 11] is a fundamental task in computer vision with various applications [31, 22, 5, 25, 21, 42, 37, 35] that involves finding the fine-grained correspondence between two images. Many image matching methods consist of two stages. First, a keypoint detector extracts collections of keypoints, which are pairs of positions and the local feature descriptors. Then, feature matchers find a match between them.
Keypoint detectors.
Classically, keypoint detection was down using hand-crafted local features such as SIFT [18] and ORB [29]. Recently, learning-based methods have been developed [7, 12, 10, 40]. Particularly, SuperPoint [7] and DISK [38] are two popular examples. The former is based on convolutional neural networks, and the latter is based on reinforcement learning. There is also a fully differentiable and lightweight method with a minimal model configuration [10]. Unsupervised keypoint detection based on a text-to-image diffusion model [12] also exists. In [40], a flexible technique, which can be combined with other keypoint detectors, is proposed to extract keypoints for robustly performing a matching under illumination changes and repeating texture patterns.
Feature matchers.
Feature matchers [43] find a fine-grained correspondence between images. The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers. The former runs faster, while the latter gives dense and high accuracy. The focus of our study lies in sparse matchers. A seminal sparse matcher, SuperGlue [30], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [39] with optimal transport [24]. The state-of-the-art sparse matcher, LightGlue [17], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously. GlueStick [23] utilizes line segments for structural features. In contrast to sparse matchers, dense matchers [36, 26, 48, 14, 8, 46, 4, 45] peruse matching accuracy rather than speed. Dense matchers input an image and infer matching using all pixels as keypoints. Therefore, it does not require a keypoint detector so it is also called a detector-free model. LoFTR [36] proposed the first transformer-based method for dense matching, which achieves highly dense matching. Recent studies such as ASpanFormer [4], which controls the attention span, and ASTR [45], which focuses on the role of each pixel.
The aforementioned methods are all designed for image matching and do not perform well in our new task, where objects to be matched do not necessarily appear the same and/or images are corrupted (cf. Fig. 6). There are a few studies that aim at robust matching in particular cases. For example, DarkFeat [11] address matching in low-light conditions, and SAM [19] peruses robustness against large viewpoint change and illumination change. As for such image corruptions, we consider fifteen types of common corruptions [13], including large brightness change and contrast change.
3 Non-Identical Object Matching
This study introduces non-identical object matching, matching between unnecessarily identical objects in two images. For example, a dog and a cat are not identical, but still have similar structures such as two eyes, four legs, and one tail. The same mass-produced products, such as cars of the same model, have the same appearance but still not identical (e.g., owned by different people). Thus, they can appear in images with largely different backgrounds and lighting conditions.
3.1 Task formulation
In the followings, we present a more formal problem setup based on a standard matching pipeline. Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
Problem 1 (Non-identical object matching)
Let ${\mathcal{X}}$ be the image domain. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ pairs of keypoint $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and its descriptor $\bm{d}_{i}^{A}\in[0,1]^{d}$ of image $\bm{x}_{A}$. We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly. Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function. The non-identical object matching can be formalized as follows.
$\min\quad$ (${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
$\mathrm{s.t.}\quad$ $m_{kl}\in[0,1],\forall k\in\{1,\ldots,n_{A}\},l\in\{1,\ldots,n_{B}\}$ (2)
$\sum_{l=1}^{n_{A}}m_{kl}\leq 1,\forall k\in\{1,\ldots,n_{A}\}$ (3)
$\sum_{k=1}^{n_{B}}m_{kl}\leq 1,\forall l\in\{1,\ldots,n_{B}\}.$ (4)
The matching coefficient $m_{kl}=1$ indicates that the $k$-th keypoint in image $\bm{x}_{A}$ and the $l$-th keypoint in image $\bm{x}_{B}$ is matched and otherwise $m_{kl}=0$. Each keypoint in one image has at most one matched point in the other image. This condition is implemented by the last two constraints. In the image matching task, where an identical object is captured from different angles in two images, the matching loss (${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
Our task, non-identical object matching, covers a broader class of matching. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{i}$ can be characterized by several attributes $(o_{i},y_{i},P_{i},D_{i})$ for $i\in\{1,2\}$. If $o_{A}=o_{B}$, then the objects to be matched are identical. If $o_{A}\equiv o_{B}$, they are the same in appearance but not necessarily identical (e.g., two cars of the same model). If $y_{A}=y_{B}$, they belong to the same class (e.g., the car class). If $P_{A}=P_{B}$, then they are captured from the same camera pose. If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
The classical image matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$. In words, it considers identical objects captured from different camera poses in the same domain. In contrast, non-identical object matching encompasses $y_{A}\neq y_{B}$ (consequently, $o_{A}\neq o_{B}$ and $o_{A}\not\equiv o_{B}$), $P_{A}\neq P_{B}$, and $D_{A}\neq D_{B}$. Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense. Below, we highlight two important special cases. For both cases, one may assume $P_{A}\neq P_{B}$.
Figure 2: Failure cases of non-identical object matching by SuperPoint [7] as a keypoint detector and LightGlue [17] as the state-of-the-art matcher. (a) The images are matched with backgrounds that have no correspondence. (b) The texture of the bear and the background are matched. (c) The shape of the corruption called frost is matched, because the matcher does not know the existence of the object. (d) There is a large difference in the matches between the right and left eyes. This is because the matcher does not consider the semantics of the object.
Class discrepancy $(y_{A}\neq y_{B})$.
Four-legged animals, such as tigers and cats, are objects of different classes, but have similar structures. They both have two eyes, four legs, and one tail. We humans can find a part-by-part matching between them, even when there are differences in color, texture, and shape. However, it is not trivial how one can design a matching methods to find a fine-graind correspondence while being robust against the difference in details. If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
Figure 3: Pipeline of the matching. The keypoint detention and feature matching are done by off-the-self models. The proposed weighting module computes the heatmap scores of objects and weights the descriptors with these semantic information.
Domain shift $(o_{A}\equiv o_{B},D_{A}\neq D_{B})$.
We humans can know the flower in the painting on the well depicts that in the garden. This is a robust fine-graind matching of an identical object in different domains. Typical domain shifts include drawings v.s. photos and clean images v.s. corrupted images. For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting. Compared to photos, drawings have less colors, textures, and shadows. Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust. If robust matching between different domains is realized, for example, one can design a system that guides a worker to assemble parts from the first-person video by comparing the parts with those in a blueprint.
3.2 Technical challenges
Generalization from image matching to non-identical object matching introduces ambiguities and perturbations into matching. First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework. There is not clear measure of a good matching that can be used for computing a loss to minimize. Second, domain shift by common corruptions or image style significantly changes the low-level features. To our knowledge, existing matching methods only use low-level features. such as colors, edges, and textures. As a consequence, these methods are susceptible to the domain shift, as shown in Fig. 2. The first challenge suggests that it would be better to somehow robustifies pretrained image matching models instead of training a model from scratch. The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.
4 Proposed Method
In this section, we present a simple plug-and-play weighting module that can turn various sparse matching models to work for non-identical object matching. Most sparse matching models, such as [30, 17, 3, 23], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher. Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, keypoint detectors output collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$. The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively. Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
4.1 Weighting module
The proposed weighting module uses an object detector and visual explanation model. In this study, we used YOLOv7 [41] and Grad-CAM [32]. Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
First, the object detector is applied to $\bm{x}$ and detects the objects with their classes. For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores, or a heatmap $H_{k}$. Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$. Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$. Finally, the weighted descriptor $\tilde{d}_{i}$ is computed by
$\widetilde{\bm{d}}_{i}=\frac{1+H_{\bm{p}_{i}}}{\underset{j\in\{1,\cdots,n\}}{\max}(1+H_{\bm{p}_{j}})}\cdot\bm{d}_{i}\in[0.5,1]^{d}.$ (5)
We empirically found that the addition of base value 1 and the normalization by the maximum value is important. For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarce of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$. Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched. If one instead uses $\widetilde{\bm{d}}_{i}=(1+H_{\bm{p}_{i}})\bm{d}_{i}\in[1,2]^{d}$, this also affects negatively. We consider that this is because of the softmax operations used in the subsequent feature matching. We discuss the consequence of the weighting of descriptors next.
4.2 Effect of weights in feature matcher
We now illustrate how our weighting module affects the subsequent feature matching. We take LightGlue [17] as an example, but a similar discussion holds for other feature matchers. For simplicity, let $\alpha_{i}$ be the weight in Eq. 5, i.e., $\widetilde{\bm{d}}_{i}=\alpha_{i}\bm{d}_{i}$. The feature matcher takes as input the keypoints $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$. The vectors $\widetilde{\bm{d}}_{i}^{A}=\alpha_{i}^{A}\bm{d}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}=\alpha_{i}^{B}\bm{d}_{i}^{B}$ are the $i$-th weighted descriptors of image $\bm{x}_{A},\bm{x}_{B}$, respectively. Below, symbols with and without $\widetilde{\,\cdot\,}$ present those with and without weighting.
In the feature matcher, the descriptors are first converted to key and query vectors using matrices $W_{q},W_{k}\in{\mathbb{R}}^{d\times d}$.
$\widetilde{\bm{q}}_{i}^{I}$ $=W_{q}\widetilde{\bm{d}}_{i}^{I}=\alpha_{i}W_{q}\bm{d}_{i}^{I}=\alpha_{i}\bm{q}_{i}^{I}$ (6)
$\widetilde{\bm{k}}_{i}^{I}$ $=W_{k}\widetilde{\bm{d}}_{i}^{I}=\alpha_{i}W_{k}\bm{d}_{i}^{I}=\alpha_{i}\bm{k}_{i}^{I},$ (7)
where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as follows. [^1] To be precise, this score should be considered as the pre-activation score because a softmax operation will be subsequently applied. The terminology of this paper follows the related studies, e.g., [17].
$\widetilde{a}_{ij}^{I}$ $=\widetilde{\bm{q}}_{i}^{I\top}\mathcal{R}\left(\bm{p}_{j}^{I}-\bm{p}_{i}^{I}\right)\widetilde{\bm{k}}_{j}^{I}$ (8)
$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot\quantity(\bm{q}_{i}^{I\top}\mathcal{R}\left(\bm{p}_{j}^{I}-\bm{p}_{i}^{I}\right)\bm{k}_{j}^{I})$ (9)
$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot a_{ij}^{I}.$ (10)
where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [33]. Therefore, our weighting by heatmap scores leads to weighting of attention scores. The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher. Roughly speaking, nothing that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0$ for between those on the background. This is reasonable because we want nice feature representations for object keypoints for nice matching. Nevertheless, such nice representations should be obtained by taking into account the background conditions, e.g., to compensate for the overall lighting conditions. If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
Similarly, our weighting affects the cross attention scores. The cross-attention score $a_{ij}^{AB}$ between the $i$-th keypoint of image $\bm{x}_{A}$ and the $j$-th keypoint of image $\bm{x}_{B}$ is computed from the key vectors.
$\widetilde{a}_{ij}^{AB}=\widetilde{\bm{k}}_{i}^{A\top}\widetilde{\bm{k}}_{j}^{B}=(\alpha^{A}_{i}\alpha^{B}_{j})\cdot a_{ij}^{AB}.$ (11)
As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ are generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
5 Experiments
We now demonstrate the effectiveness of our method in non-identical object matching tasks. Particularly, we first present qualitative results on various pairs of non-identical (but similar) objects. We then quantitatively evaluate the proposed methods on classical image matching tasks but under common corruptions.
Setup.
In the sparse matching framework, the image matching task is performed by a keypoint detector and a feature matcher. For the former, we adopted pre-trained SuperPoint [7], extracting 2,048 keypoints. For the latter, the extracted keypoints are matched using either of pretrained SGMNet [3], GlueStick [23], and the state-of-the-art method, LightGlue [17], with and without the proposed weighting module. We additionally tested a detector-free dense matcher, LoFTR [36]. We downloaded the pretrained models from their official repositories.222 SGMNet: https://github.com/vdvchen/SGMNet
GlueStick: https://github.com/cvg/GlueStick
LightGlue: https://github.com/cvg/LightGlue
LoFTR: https://github.com/zju3dv/LoFTR
All the experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
5.1 Matching Non-Identical Objects
In this experiment, we address a matching of non-identical objects. It is worth noting that this task can only be qualitatively evaluated because there is no ground truth matching between non-identical objects.
Datasets.
We collected various images with objects from the COCO [16] and ImageNet [6] datasets. To prepare the drawing version of several images, we used the DreamStudio [^3] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [28].
Robust matching across non-identical objects.
Figure 4: Typical categories of non-identical object matching. (a) A combination of SuperPoint [7] and LightGlue [17] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching.
We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) different class (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) another domain (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$). For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$). Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in the supplementary material. Figure 4 shows the results of the matching results with and without the proposed method in the four cases. As can be seen in Fig. 4(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals. In contrast, Fig. 4(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings. Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched rear wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles. Notably, structural features, such as eyes and noses, are successfully matched between different classes. We consider that the heatmaps of these animals gave similar importance to the eyes or mouth, thus increasing the number of matches between the same parts of the body. In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
Different matchers.
Figure 5 presents the results with different feature matchers, showing that the proposed method boosts all of them to match corresponding parts of the cheetah and Siberian husky.
Figure 5: The results with different feature matchers [3, 23, 17]. SuperPoint [7] is used for the keypoint detection. The introduction of the proposed module increases the matching of eyes and noses between the cheetah and Siberian husky and decreases mismatching in the background.
5.2 Robust Image Matching
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
Gaussian Noise 43.09 53.27 45.67 52.24 34.97
Shot Noise 43.41 53.82 46.10 51.72 40.00
Impulse Noise 44.98 50.11 44.91 50.05 37.73
Defocus Blur 32.69 48.35 45.58 49.23 57.22
Frosted Glass Blur 33.37 47.88 47.40 49.94 54.48
Motion Blur 42.12 53.63 44.76 51.84 55.60
Zoom Blur 24.40 31.21 22.87 30.97 23.08
Snow 31.51 30.33 25.26 23.21 31.85
Frost 32.24 31.71 26.39 24.22 19.23
Fog 70.99 73.49 71.19 73.89 67.10
Brightness 75.48 75.08 73.03 70.84 75.45
Contrast 39.50 38.47 39.47 42.61 53.22
Elastic Transform 54.78 66.21 60.37 64.36 58.56
Pixelate 67.94 68.24 65.01 64.81 74.77
Jpeg Compression 29.54 36.97 39.66 44.70 59.05
Average 44.40 50.58 46.51 49.64 49.49
Time per pair [msec.] 46.67 68.43 105.72 128.18 312.54
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 80.61 78.42 78.04 75.08 80.93
Gaussian Noise 27.54 41.01 41.93 44.98 33.02
Shot Noise 32.10 42.35 41.75 43.82 36.53
Impulse Noise 35.95 43.67 40.97 42.91 34.92
Defocus Blur 18.25 32.13 23.16 30.10 49.40
Frosted Glass Blur 34.25 44.80 42.96 46.21 54.53
Motion Blur 50.07 53.10 44.27 52.69 52.81
Zoom Blur 34.67 40.04 31.18 37.11 31.09
Snow 56.24 58.15 47.22 48.53 49.15
Frost 62.20 64.87 54.34 57.26 43.17
Fog 76.96 78.40 75.18 74.87 74.06
Brightness 77.14 76.98 76.78 75.06 78.41
Contrast 43.09 45.17 40.44 43.53 39.66
Elastic Transform 64.93 68.34 68.02 69.92 68.95
Pixelate 66.80 68.43 67.16 65.46 76.21
Jpeg Compression 47.41 53.82 60.39 60.12 68.14
Average 48.51 54.08 50.38 52.84 52.67
We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions. To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions. We address two cases: (i) both input images are corrupted and (ii) only one of them is corrupted. The former shows the robustness against corruptions, and the latter shows the robustness against domain shift (i.e., clean image domain v.s. corrupted image domain).
Dataset.
We use image pairs from the MegaDepth-1500 [36] test set as the basis for various experiments. MegaDepth-1500 is a set of 1,500 image pairs from the MegaDepth [15] dataset for two popular phototourism destinations. MegaDepth consists of 3D reconstructed data from one million images, its camera poses and depth maps are first computed from COLMAP [31] and then refined as ground truth. Common Corruptions [13] simulates the 15 types of corruptions that occur when images are captured. The severity level of corruption is set to 5, which is the most challenging condition.
Metric.
[…]
Robustness against common corruptions.
Table 1 shows the AUC of matching when input images are both corrupted. The results of the dense matching method (i.e., LoFTR) is given for reference. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight increase in the runtime. The average AUC even exceeds that of LoFTR. The matching methods benefit from the proposed method particularly when images are corrupted by noises and blurs. We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method. These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them. See Section B.1 in the supplementary materials for more detailed comparisons.
Robustness against environmental changes.
As the objects to be matched are non-identical, they can appear in largely different environments. Table 2 shows the AUC when only either of the input images is corrupted and the other is clean. As in Tab. 1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case. Interestingly, the results are better for almost all weather categories. For brightness, it is the only one of all categories that shows a decrease in accuracy, but the drop is slight. Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions. Heatmap weighting is considered to structurally complement the areas hidden by occlusion. The proposed method makes existing matchers robust to environmental changes in general, with a slight precision reduction in clean images. See Section B.2 in the supplementary material for more detailed comparisons.
Figure 6: Failure cases. (a) High heatmap scores are given to only a small region of objects to be matched. (b) The structures of objects are clearly different.
5.3 Limitations
We here present failure cases. Figure 6 shows typical cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences. In the former case, the importance scores take large only on particular parts of the objects. Further, the large importance scores on the irrelevant object disrupt the matching. This may be resolved by using a better visual explanation model [47, 32, 2, 1, 20, 34]. For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched. To address these cases, one may introduce recent vision and language models to complement the visual information with text features.
6 Conclusion
In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical as in prior studies. Because of the non-identicality, two objects may be captured in different environments and domains. Thus, this task also encompasses robust matching under image corruptions and domain shifts. The proposed method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects. Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP22H03658 and JP22K17962. We thank Satomi Tanaka (Ricoh Company, Ltd.) for her fruitful comments and financial support from Ricoh Company, Ltd..
Supplementary Material
A More matching non-identical objects
In Sec. 5.1 of the main text, we evaluated non-identical object matching by inserting the proposed method between SuperPoint [7] and LightGlue [17]. The matching results for various image pair categories show that the proposed method enhances existing matchers and provides dense and consistent correspondences. To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$).
Setup.
We collected various images with objects from the COCO [16] and ImageNet [6] datasets. To prepare drawing or sketch versions of several images, we used the DreamStudio [^4] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [28]. For details of the experiment, refer to Sec. 5.1.
A.1 Success cases
Figure 7 shows the matching results for the same class case. As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one. The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching. Even in the case of class discrepancy, as shown in Fig. 8, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals. Figure 9 shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings). These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
A.2 Failure cases
There are some cases of failure in non-identical object matching, as shown in Fig. 10. It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller), or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object. Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
Figure 7: Matching results in the same class case. Image pairs show objects of the same class in ImageNet [6] or COCO [16] datasets. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching, and performs dense and consistent correspondences.
Figure 8: Matching results in the class discrepancy case. Image pairs show objects with class discrepancies in ImageNet [6] or COCO [16] datasets. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence. (b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
Figure 9: Matching results in the domain shift case. Image pairs for matching are ImageNet [6] or COCO [16] datasets with common corruptions [13], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [28]. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] finds correct matching, but also finds mismatches with many non-correspondence areas. (b) The proposed method reduces these mismatches and further improves to denser matching between objects.
Figure 10: Failure cases of non-identical object matching. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] fails in matching. (b) The proposed method improves several matchings but still leaves many corresponding areas unmatched.
B Robust image matching details
[…]
B.1 Robustness against common corruptions.
[…]
B.2 Robustness against environmental changes.
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88
Gaussian Noise 27.44 37.28 28.64 35.11 20.07
Shot Noise 27.07 39.79 27.81 33.42 24.27
Impulse Noise 28.70 37.25 29.10 34.44 22.27
Defocus Blur 18.58 24.51 30.12 33.78 43.27
Frosted Glass Blur 17.24 27.61 29.16 31.75 39.21
Motion Blur 27.08 34.85 27.98 35.46 40.68
Zoom Blur 12.50 20.23 12.64 20.72 12.01
Snow 18.35 12.56 12.46 10.49 18.94
Frost 20.86 18.75 17.64 15.48 11.15
Fog 56.32 62.62 55.14 57.24 53.05
Brightness 61.10 63.95 58.75 56.51 61.88
Contrast 24.33 23.10 24.31 27.45 38.57
Elastic Transform 36.76 47.02 41.53 45.52 41.27
Pixelate 51.60 60.66 47.60 47.32 62.55
Jpeg Compression 16.60 27.73 22.35 27.64 43.49
Average 29.64 35.86 31.02 34.16 35.51
[…]
Gaussian Noise 14.70 24.44 14.63 21.53 9.85
Shot Noise 13.71 23.44 13.44 18.06 12.21
Impulse Noise 15.65 21.25 16.90 20.64 10.92
Defocus Blur 8.22 11.69 16.59 20.64 27.77
Frosted Glass Blur 7.17 18.09 14.58 16.12 24.59
Motion Blur 14.72 20.72 15.26 22.34 25.37
Zoom Blur 5.38 8.82 6.02 14.16 5.17
Snow 9.48 8.32 5.92 3.87 9.91
Frost 11.92 11.29 9.92 7.72 6.09
Fog 39.01 43.43 35.88 37.58 36.94
Brightness 43.60 47.98 41.86 39.60 45.13
Contrast 11.69 13.93 12.75 15.29 23.13
Elastic Transform 20.43 29.64 24.29 27.28 24.40
Pixelate 32.97 41.82 29.86 29.66 46.40
Jpeg Compression 8.00 15.94 9.71 14.79 27.77
Average 17.11 22.72 17.84 20.62 22.38
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88
Gaussian Noise 16.63 28.34 28.57 32.42 24.17
Shot Noise 19.97 32.50 26.23 29.92 26.58
Impulse Noise 23.56 34.74 33.12 36.26 25.67
Defocus Blur 8.29 22.43 12.03 18.17 36.85
Frosted Glass Blur 19.00 32.78 26.39 29.66 40.40
Motion Blur 33.29 39.69 27.88 38.30 39.41
Zoom Blur 20.38 26.23 16.76 25.67 21.60
Snow 39.74 41.07 30.74 33.00 35.91
Frost 47.61 51.78 38.80 41.62 32.97
Fog 62.39 64.08 61.11 60.80 60.74
Brightness 63.40 61.75 62.69 60.04 65.64
Contrast 27.81 33.12 25.86 28.95 30.27
Elastic Transform 47.62 56.66 52.17 55.08 53.61
Pixelate 50.14 56.68 50.33 48.63 63.14
Jpeg Compression 29.92 40.31 40.93 40.61 53.58
Average 33.98 41.48 35.57 38.61 40.70
[…]
Gaussian Noise 8.42 20.10 16.51 20.52 16.68
Shot Noise 10.30 20.52 16.14 19.31 18.14
Impulse Noise 12.88 24.11 21.41 25.35 17.75
Defocus Blur 3.22 17.92 5.95 15.89 24.73
Frosted Glass Blur 8.56 22.42 15.54 18.79 26.87
Motion Blur 18.95 26.22 13.94 24.36 25.92
Zoom Blur 9.81 19.27 8.95 17.68 15.03
Snow 22.94 26.90 17.27 20.56 23.94
Frost 31.58 37.80 24.13 27.95 23.07
Fog 44.49 45.00 42.65 41.24 44.04
Brightness 46.14 45.24 43.91 41.63 48.85
Contrast 14.98 17.15 14.46 17.53 21.09
Elastic Transform 29.34 39.29 32.27 35.11 36.41
Pixelate 31.51 38.02 31.82 30.10 46.26
Jpeg Compression 15.86 27.55 22.24 21.87 37.35
Average 20.60 28.50 21.81 25.19 28.41

### v3

@@
 Abstract
-Not identical but similar objects are everywhere in the world.
-Examples include four-legged animals such as dogs and cats, cars of different models, akin flowers in various colors, and countless others.
-In this study, we address a novel task of matching such non-identical objects.
-We propose a simple weighting scheme of descriptors that enhance various sparse image matching methods, which are originally designed for matching identical objects captured from different perspectives, and achieve semantically robust matching.
-The experiments show successful matching between non-identical objects in various cases including domain shift.
-Further, we present a first evaluation of the robustness of the image matching methods under common corruptions, which is a sort of domain shift, and the proposed method improves the matching in this case as well.
+Not identical but similar objects are ubiquitous in our world, ranging from four-legged animals such as dogs and cats to cars of different models and flowers of various colors.
+This study addresses a novel task of matching such non-identical objects at the pixel level.
+We propose a weighting scheme of descriptors that incorporates semantic information from object detectors into existing sparse image matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects.
+The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo–drawing and image corruptions).
 1 Introduction
-Figure 1: Non-identical object matching is achieved densely and consistently by adding our method between SuperPoint [7] and LightGlue [17].
-Object correspondences can be found even when the design is different as in (a), the classes are discrepant as in (b), and there are several similar objects as in (c).
-Objects are robustly matched to each other even with domain shifts in pictures and photos as in (d) and corruption as in (e) and (f).
-We humans are able to find fine-grained visual correspondences between two objects even when they look similar but not identical or when one is sketched and another appears in a photo.
-Thanks to this ability, for example, we can identify a cat in gloomy backstreet from a photo taken in a bright room or assemble components to build a shelf based on its perspective drawing.
-In computer vision words, this needs fine-grained matching between feature points of two objects under image corruptions, class shift, and domain shift.
-The classical image matching task [30, 36, 17, 11] only addresses a special case; objects to be matched are assumed to be identical and captured from different perspectives.
-Recent Transformer-based dense matching methods show a great performance on such a matching task [36, 26, 48, 14, 8, 46, 4, 45], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [30, 3, 44, 27, 23, 11, 19].
-However, they do not perform well once the identicality assumption is not violated and/or objects are captured in harsh environments or different domains (cf.
-Fig.
+We humans are able to find fine-grained and robust visual correspondences between two objects even when they are not necessarily identical or when they are captured in different situations.
+Thanks to this ability, for example, we can find out our missing cats in a gloomy backstreet by remembering them basking in the sun or assembling components to build a shelf based on their perspective drawing.
+In computer vision words, this needs fine-grained matching of feature points between two objects even with class discrepancy and domain shift (e.g., image corruptions).
+The abovementioned matching has been partially addressed in the literature.
+In the image matching [49, 56, 27, 17], objects to be matched are assumed identical and captured from different perspectives.
+This assumption also enables supervised training using dense correspondence obtained from 3D reconstruction from image pairs.
+Recently, Transformer-based dense matching methods have shown great performance [56, 45, 74, 23, 11, 72, 5, 67], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [49, 66, 46, 38, 17, 30, 22].
+In contrast, semantic correspondence models [70, 32, 7] focus on matching semantically similar parts between objects (e.g., tires of two bikes).
+They are trained on datasets with sparse part-to-part correspondence annotation and thus do not offer fine-grained matching.
+Figure 1: Non-identical object matching by our plug-and-play module incorporated with the sparse image matching pipeline (here, SuperPoint [10] for keypoint detection and LightGlue [27] for matching).
+(a) design differences, (b) class discrepancy, (c) similar objects, and domain shifts such as (d) drawing v.s.
+photo and (e, f) different image corruptions.
+In this study, we generalize image matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig.
+1).
+For example, the object pair can be street clocks of different designs or four-legged animals of different species.
+Further, the non-identicality naturally requires robustness, e.g., against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s.
+drawing) because two objects may be captured in different environments and domains.
+A crucial challenge of achieving non-identical object matching is that we cannot straightforwardly resort to the supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
+We propose a simple plug-and-play weighting module that extends off-the-shelf sparse matching methods to non-identical object matching.
+Given a pair of images, existing sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints.
+Our weighting module is inserted before the feature matching model and weights the descriptors using semantic information.
+Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [51] to obtain weights.
+Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in the non-identical object matching task (see Fig.
 2).
-In this study, we address a novel task of non-identical object matching, where objects to be matched are not necessarily identical.
-For example, the objects to be matched can be clocks of different designs or animals of different species.
-Further, the non-identicality naturally induces the necessity of robustness against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s.
-drawing) because two objects may be captured in different environments and domains, as shown in Fig.
-1.
-A crucial challenge of achieving non-identical object matching is that we cannot simply resort to a supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
-At least, to our knowledge, there are no such datasets.
-Interestingly, we discovered a simple plug-and-play weighting module can extend off-the-shelf sparse matching methods to the non-identical object matching.
-Namely, given a pair of images, sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints.
-Our weighting module can be inserted before the feature matching model to weights the descriptors using semantic information.
-Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [32] to obtain weights.
-Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in non-identical object matching task.
 The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features.
-Our experiments demonstrate that this simple approach significantly boosts the robustness of sparse matching methods, including X and Y, in two tasks: non-identical object matching and robust image matching.
-In the former task, we evaluate the matching results with the image pairs (a) of the same class, (b) of different classes, (c) with domain shift, and (d) of the same appearance.
-The sparse matchers alone resulted in scattered mismatching between non-corresponding areas, but with the addition of our method, fine-grained and consistent matching is achieved.
+Further, to avoid emphasizing irrelevant objects when an image contains more than one object, we propose Class Similarity Adapter, which determines the most relevant object pairs.
+We leverage the class labels from an object detector and determine object pairs based on their semantic similarity in the text embedding space.
+Our experiments demonstrate that our approach significantly boosts the robustness of sparse matching methods, including LightGlue [27] and GlueStick [38], in two tasks: non-identical object matching and robust image matching.
+In the former task, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance.
+The sparse matchers alone result in scattered mismatching between non-corresponding areas, but combined with our method, fine-grained and consistent matching is achieved.
 In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [56] dataset under common corruptions [20] in terms of the AUC score of standard relative pose estimation.
-The proposed method increases the AUC scores with various types of common corruptions, and exceeds a dense matching method LoFTR [36] in the average case.
+Our method outperforms the state-of-the-art sparse matcher under various types of image corruptions.
+The average AUC of our method even exceeds that of dense matchers, LoFTR [56] and Efficient LoFTR [64], which pursue matching quality at the cost of speed.
 To summarize, our contributions are as follows:
-  • We propose non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images even under image corruptions and/or between different domains.
-This task can be considered a generation of the classical image matching task.
+  • We tackle non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images, even under image corruptions and/or between different domains, which can be considered a generation of classical image matching and semantic correspodense.
   • We propose a simple plug-and-play weighting module that enhances the low-level features using semantic information and successfully extends various sparse matching methods to non-identical object matching.
   • We present a first evaluation of the robustness in image matching under common corruptions.
-By adding corruptions to both input images, the robustness of corruptions itself is measured.
+By adding corruptions to both input images, the robustness against corruptions itself is measured.
 By adding them to only one of the input images, the robustness against a domain shift is measured.
-For both cases, the proposed method is shown useful.
+For both cases, our method is shown useful.
 2 Related Work
+Figure 2: Failure cases of non-identical object matching using the state-of-the-art matcher, LightGlue [27], and the semantic correspondence model, GeoAware-SC [70].
+In semantic correspondence models, some points indicating object parts must be designated during inference to find correspondences, and automating this point selection is essential for feature point matching.
+Comparison of feature matching by GeoAware-SC using two methods: one that randomly extracts 50 points and the other that adaptively extracts 50 points based on the GradCAM [51] heatmap of YOLOv7 [62].
+Red boxes indicate mismatches due to background or object-like textures.
+Orange boxes indicate semantically corresponding regions, where most points align at a part level but not at a pixel level.
+For example, in (d), mismatches occur between the tip of the nose of the cheetah and the nostrils of the husky.
 Image matching [49, 56, 27, 17] is a fundamental task in computer vision with various applications [50, 35, 8, 43, 34, 63, 57, 55] that involves finding the fine-grained correspondence between two images.
-Many image matching methods consist of two stages.
-First, a keypoint detector extracts collections of keypoints, which are pairs of positions and the local feature descriptors.
-Then, feature matchers find a match between them.
+Many image matching methods have two stages.
+A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find correspondences between them.
 Keypoint detectors.
-Classically, keypoint detection was down using hand-crafted local features such as SIFT [18] and ORB [29].
-Recently, learning-based methods have been developed [7, 12, 10, 40].
+Classically, keypoint detection was done using hand-crafted local features such as SIFT [28] and ORB [48].
+Recently, learning-based methods have been developed [10, 18, 14, 61, 41, 12].
 Particularly, SuperPoint [10] and DISK [59] are two popular examples.
@@
 Unsupervised keypoint detection based on a text-to-image diffusion model [18] also exists.
-In [40], a flexible technique, which can be combined with other keypoint detectors, is proposed to extract keypoints for robustly performing a matching under illumination changes and repeating texture patterns.
 Feature matchers.
-Feature matchers [43] find a fine-grained correspondence between images.
+Feature matchers [65, 68] find a fine-grained correspondence between images.
 The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers.
@@
 Dense matchers input an image and infer matching using all pixels as keypoints.
-Therefore, it does not require a keypoint detector so it is also called a detector-free model.
+Therefore, it does not require a keypoint detector, so it is also called a detector-free model.
 LoFTR [56] proposed the first transformer-based method for dense matching, which achieves highly dense matching.
-Recent studies such as ASpanFormer [4], which controls the attention span, and ASTR [45], which focuses on the role of each pixel.
-The aforementioned methods are all designed for image matching and do not perform well in our new task, where objects to be matched do not necessarily appear the same and/or images are corrupted (cf.
-Fig.
-6).
-There are a few studies that aim at robust matching in particular cases.
-For example, DarkFeat [11] address matching in low-light conditions, and SAM [19] peruses robustness against large viewpoint change and illumination change.
-As for such image corruptions, we consider fifteen types of common corruptions [13], including large brightness change and contrast change.
+Recent studies [36, 71, 29] such as Efficient LoFTR [64], enabling fast and accurate matching by token aggregation, ASpanFormer [5], which controls the attention span, and ASTR [67], which focuses on the role of each pixel.
+Semantic correspondence models.
+Semantic Correspondence [15, 16, 33, 6, 2] is a task of identifying a semantically similar point in the target image from the given point in the source image.
+Specifically, it enables the correspondence of parts, such as the mouth, nose, or right eye of dogs between the source and target images.
+The training of a model is conducted using a specialized dataset, such as SPair-71k [33] and PF-PASCAL [15], which employ sparsely annotated part-level correspondences as supervisory data.
+Recently, GeoAware-SC [70] has been proposed as a model that leverages pre-trained features from two foundational models, Stable Diffusion [47] and DINOv2 [37], integrating both geometric and semantic information.
+This model significantly improves accuracy on major benchmarks for semantic correspondence, including the SPair-71k [33] and PF-PASCAL [15] datasets.
+In recent years, various models [39, 52, 69, 58, 21] have been studied.
+For example, CATs++ [7], which combines the local feature aggregation capacity of CNNs with the broad contextual comprehension of transformers; a model [32] employing feature maps from the reverse diffusion process in diffusion models; and unsupervised models [19] that learn semantic correspondence without requiring annotated data.
+While feature matching and semantic correspondence are also the tasks of finding correspondences, our task (i.e., fine-grained matching between non-identical objects) is more challenging.
+First, it is hard to be addressed by the supervised learning framework because (i) manual annotation, as in semantic correspondence datasets is laborious for the pixel-level correspondence, and (ii) automatic annotation by 3D reconstruction, as in feature matching, which assumes an identical object viewed from different perspectives, cannot be done.
+Second, because non-identical objects generally appear with different lighting, image corruptions, or even image domains, the matching method should be designed to be robust to such diversity.
+As shown in Fig.
+2, state-of-the-art feature matching methods and semantic correspondence models are susceptible to such variations.
+Several studies addressed robust matching with specific conditions, such as DarkFeat [17] for low-light and SAM [30] for viewpoint and illumination changes.
+We consider broader changes, including fifteen types of common corruptions [20] and several image domains (e.g., photos and sketches).
 3 Non-Identical Object Matching
-This study introduces non-identical object matching, matching between unnecessarily identical objects in two images.
-For example, a dog and a cat are not identical, but still have similar structures such as two eyes, four legs, and one tail.
-The same mass-produced products, such as cars of the same model, have the same appearance but still not identical (e.g., owned by different people).
-Thus, they can appear in images with largely different backgrounds and lighting conditions.
+We introduce non-identical object matching, matching between unnecessarily identical objects in two images.
+For example, a dog and a cat are not identical but still have similar structures, such as two eyes, four legs, and one tail.
+The same mass-produced products, such as cars of the same model, have the same appearance but are still not identical (e.g., owned by different people) and can appear in images with largely different backgrounds and lighting conditions.
 3.1 Task formulation
-In the followings, we present a more formal problem setup based on a standard matching pipeline.
+In the following, we present a more formal problem setup based on a standard matching pipeline.
 Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
@@
 Let $\mathcal{X}$} be the image domain.
-Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ pairs of keypoint $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and its descriptor $\bm{d}_{i}^{A}\in[0,1]^{d}$ of image $\bm{x}_{A}$.
+Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are the position and the descriptor of $i$-th keypoint, respectively.
 We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly.
@@
 The non-identical object matching can be formalized as follows.
-$\min\quad$ (${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
+$\min\quad$ ${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
 $\mathrm{s.t.}\quad$ $m_{kl}\in[0,1],\forall k\in\{1,\ldots,n_{A}\},l\in\{1,\ldots,n_{B}\}$ (2)
@@
 This condition is implemented by the last two constraints.
-In the image matching task, where an identical object is captured from different angles in two images, the matching loss (${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
+In the image matching task, where an identical object is captured from different angles in two images, the matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
 Our task, non-identical object matching, covers a broader class of matching.
-Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{i}$ can be characterized by several attributes $(o_{i},y_{i},P_{i},D_{i})$ for $i\in\{1,2\}$.
+Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{I}$ can be characterized by several attributes $(o_{I},y_{I},P_{I},D_{I})$ for $I\in\{A,B\}$.
 If $o_{A}=o_{B}$, then the objects to be matched are identical.
@@
 Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense.
-Below, we highlight two important special cases.
-For both cases, one may assume $P_{A}\neq P_{B}$.
-Figure 2: Failure cases of non-identical object matching by SuperPoint [7] as a keypoint detector and LightGlue [17] as the state-of-the-art matcher.
-(a) The images are matched with backgrounds that have no correspondence.
-(b) The texture of the bear and the background are matched.
-(c) The shape of the corruption called frost is matched, because the matcher does not know the existence of the object.
-(d) There is a large difference in the matches between the right and left eyes.
-This is because the matcher does not consider the semantics of the object.
+Below, we highlight two important special cases, both assuming $P_{A}\neq P_{B}$.
+Figure 3: Pipeline of the matching.
+The keypoint detection and feature matching are done by off-the-shelf models.
+The proposed weighting module computes the heatmap scores of objects and weights the descriptors with this semantic information.
 Class discrepancy $(y_{A}\neq y_{B})$.
-Four-legged animals, such as tigers and cats, are objects of different classes, but have similar structures.
+Four-legged animals, such as tigers and cats, are objects of different classes but have similar structures.
 They both have two eyes, four legs, and one tail.
-We humans can find a part-by-part matching between them, even when there are differences in color, texture, and shape.
-However, it is not trivial how one can design a matching methods to find a fine-graind correspondence while being robust against the difference in details.
+We humans can find a part-by-part matching between them, even with differences in color, texture, and shape.
+However, it is not trivial to design matching methods to find fine-grained correspondence while being robust to detail differences.
 If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
-Figure 3: Pipeline of the matching.
-The keypoint detention and feature matching are done by off-the-self models.
-The proposed weighting module computes the heatmap scores of objects and weights the descriptors with these semantic information.
 Domain shift $(o_{A}\equiv o_{B},D_{A}\neq D_{B})$.
-We humans can know the flower in the painting on the well depicts that in the garden.
-This is a robust fine-graind matching of an identical object in different domains.
+We humans can know the flower in the painting on the wall depicts that in the garden.
+This is a robust fine-grained matching of an identical object in different domains.
 Typical domain shifts include drawings v.s.
@@
 For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting.
-Compared to photos, drawings have less colors, textures, and shadows.
+Compared to photos, drawings have fewer colors, textures, and shadows.
 Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust.
@@
 First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework.
-There is not clear measure of a good matching that can be used for computing a loss to minimize.
-Second, domain shift by common corruptions or image style significantly changes the low-level features.
-To our knowledge, existing matching methods only use low-level features.
-such as colors, edges, and textures.
+There is no clear measure of a good matching that can be used for computing a loss to minimize.
+Second, domain shifts by common corruptions or image styles significantly change the low-level features.
+To our knowledge, existing matching methods only use low-level features such as colors, edges, and textures.
 As a consequence, these methods are susceptible to the domain shift, as shown in Fig.
 2.
-The first challenge suggests that it would be better to somehow robustifies pretrained image matching models instead of training a model from scratch.
+The first challenge suggests that it would be better to somehow robustify pretrained image matching models instead of training a model from scratch.
 The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.
 4 Proposed Method
-In this section, we present a simple plug-and-play weighting module that can turn various sparse matching models to work for non-identical object matching.
-Most sparse matching models, such as [30, 17, 3, 23], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher.
-Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, keypoint detectors output collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$.
+In this section, we present a simple plug-and-play weighting module that can turn various sparse matchers to work for non-identical object matching.
+Most sparse matchers, such as [49, 27, 4, 38], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher.
+Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, The keypoint detector outputs collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$.
 The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively.
@@
 The proposed weighting module uses an object detector and visual explanation model.
-In this study, we used YOLOv7 [41] and Grad-CAM [32].
+This study used YOLOv7 [62] and Grad-CAM [51].
 Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
 First, the object detector is applied to $\bm{x}$ and detects the objects with their classes.
-For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores, or a heatmap $H_{k}$.
+For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores or a heatmap $H_{k}$.
 Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$.
 Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$.
-Finally, the weighted descriptor $\tilde{d}_{i}$ is computed by
+Finally, the weighted descriptor $\tilde{\bm{d}}_{i}$ is computed by
 $\widetilde{\bm{d}}_{i}=\frac{1+H_{\bm{p}_{i}}}{\underset{j\in\{1,\cdots,n\}}{\max}(1+H_{\bm{p}_{j}})}\cdot\bm{d}_{i}\in[0.5,1]^{d}.$ (5)
 We empirically found that the addition of base value 1 and the normalization by the maximum value is important.
-For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarce of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$.
+For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarcity of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$.
 Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched.
@@
 5, i.e., $\widetilde{\bm{d}}_{i}=\alpha_{i}\bm{d}_{i}$.
-The feature matcher takes as input the keypoints $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
+The matcher takes as input the keypoints $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
 The vectors $\widetilde{\bm{d}}_{i}^{A}=\alpha_{i}^{A}\bm{d}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}=\alpha_{i}^{B}\bm{d}_{i}^{B}$ are the $i$-th weighted descriptors of image $\bm{x}_{A},\bm{x}_{B}$, respectively.
@@
 where $I\in\{A,B\}$.
-Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as follows.
-[^1] To be precise, this score should be considered as the pre-activation score because a softmax operation will be subsequently applied.
+Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as [^1] To be precise, this score should be considered as the pre-activation score, and the softmax operation will be subsequently applied.
 The terminology of this paper follows the related studies, e.g., [27].
@@
 $=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot\quantity(\bm{q}_{i}^{I\top}\mathcal{R}\left(\bm{p}_{j}^{I}-\bm{p}_{i}^{I}\right)\bm{k}_{j}^{I})$ (9)
-$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot a_{ij}^{I}.$ (10)
+$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot a_{ij}^{I},$ (10)
 where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [53].
-Therefore, our weighting by heatmap scores leads to weighting of attention scores.
+Therefore, our weighting by heatmap scores leads to the weighting of attention scores.
 The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher.
@@
 If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
-Similarly, our weighting affects the cross attention scores.
-The cross-attention score $a_{ij}^{AB}$ between the $i$-th keypoint of image $\bm{x}_{A}$ and the $j$-th keypoint of image $\bm{x}_{B}$ is computed from the key vectors.
+Similarly, our weighting affects the cross-attention score $a_{ij}^{AB}$ between the $i$-th keypoint of image $\bm{x}_{A}$ and the $j$-th keypoint of image $\bm{x}_{B}$, computed from the key vectors.
 $\widetilde{a}_{ij}^{AB}=\widetilde{\bm{k}}_{i}^{A\top}\widetilde{\bm{k}}_{j}^{B}=(\alpha^{A}_{i}\alpha^{B}_{j})\cdot a_{ij}^{AB}.$ (11)
-As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ are generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
+As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ is generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
+4.3 Class Similarity Adapter
+Figure 4: Class Similarity Adapter converts class labels from an object detector into embeddings using the CLIP [42] text encoder and compares them with cosine similarity to identify semantically similar object pairs.
+The GradCAM [51] heatmap of object detector YOLOv7 [62] generated for the chosen object pair is used for our weighting method, enhancing matching with multiple objects.
+Here, we used the keypoint detector SuperPoint [10] and the feature matcher LightGlue [27].
+We propose the Class Similarity Adapter to address scenarios where multiple objects are present in an image.
+Our weighting module uses all heatmaps of objects detected by the object detector for weighting.
+However, this can needlessly emphasize irrelevant objects, leading to scattered matching from one to multiple objects.
+As shown in  Fig.
+4 (bottom right), relying solely on visual information leads to matching based on similarities in texture, color, and shape, where similar black features such as the nose of the dog and the ear of the cat, are incorrectly matched, causing correspondences to scatter across multiple animals.
+To resolve such confusion in the visual space, we again leverage semantic information given by the object detector.
+Particularly, we map the class labels of detected objects to text embedding space and determine the object pair to be matched based on the cosine similarity between embedding vectors.
+Suppose that an image $\bm{x}_{A}$ contains $n_{A}$ objects, and let $Y_{A}=\{y_{i}^{A}\}_{i=1}^{n_{A}}$ be their class labels.
+We define $Y_{B}=\{y_{i}^{B}\}_{i=1}^{n_{B}}$ similarly for the other image $\bm{x}_{B}$.
+Let $\mathcal{T}$ be a pretrained text encoder that maps a label $y$ to an embedding vector $\mathcal{T}(y)$.
+Particularly, we used the CLIP text encoder [42].
+The Class Similarity Adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
+$(y_{i}^{A},y_{j}^{B})=\underset{y^{A}\in Y_{A},~y^{B}\in Y_{B}}{\text{argmax}}\;\frac{\mathcal{T}(y^{A})^{\top}\mathcal{T}(y^{B})}{\|\mathcal{T}(y^{A})\|\|\mathcal{T}(y^{B})\|},$ (12)
+where $\|\,\cdot\,\|$ denotes the $L_{2}$ norm of vector.
+Then, the weighting module only uses the Grad-CAM heatmaps of the $i$-th object in $\bm{x}_{A}$ and the $j$-th object in $\bm{x}_{B}$.
+Based on this result, the weighting module uses the corresponding heatmaps for weighting the descriptors.
+Figure 4 shows this pipeline.
 5 Experiments
-We now demonstrate the effectiveness of our method in non-identical object matching tasks.
-Particularly, we first present qualitative results on various pairs of non-identical (but similar) objects.
-We then quantitatively evaluate the proposed methods on classical image matching tasks but under common corruptions.
+We demonstrate the effectiveness of our method in non-identical object matching tasks, presenting qualitative results on various pairs of non-identical (but similar) objects, followed by quantitative evaluation on classical image matching tasks but under common corruptions [20].
 Setup.
@@
 For the former, we adopted pre-trained SuperPoint [10], extracting 2,048 keypoints.
-For the latter, the extracted keypoints are matched using either of pretrained SGMNet [3], GlueStick [23], and the state-of-the-art method, LightGlue [17], with and without the proposed weighting module.
-We additionally tested a detector-free dense matcher, LoFTR [36].
-We downloaded the pretrained models from their official repositories.222 SGMNet: https://github.com/vdvchen/SGMNet
-GlueStick: https://github.com/cvg/GlueStick
-LightGlue: https://github.com/cvg/LightGlue
-LoFTR: https://github.com/zju3dv/LoFTR
-All the experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
+For the latter, the extracted keypoints are matched using either pretrained SGMNet [4], GlueStick [38], and the state-of-the-art method, LightGlue [27], with and without the proposed weighting module.
+We additionally tested a detector-free dense matcher, LoFTR [56] and Efficient LoFTR [64].
+We downloaded the pretrained models from their official repositories (details provided in the supplementary material).
+All experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
 5.1 Matching Non-Identical Objects
@@
 To prepare the drawing version of several images, we used the DreamStudio [^2] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [47].
-Robust matching across non-identical objects.
 Figure 5: Typical categories of non-identical object matching.
@@
 (b) The proposed method significantly improves the matching.
-We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) different class (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) another domain (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$).
+Robust matching across non-identical objects.
+We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$).
 For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$).
-Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in the supplementary material.
-Figure 4 shows the results of the matching results with and without the proposed method in the four cases.
+Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in Sec.
+A in the supplementary material.
+Figure 5 shows the matching results with and without the proposed method in the four cases.
 As can be seen in Fig.
-4(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals.
+5(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals.
 In contrast, Fig.
-4(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings.
-Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched rear wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles.
+5(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings.
+Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched front wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles.
 Notably, structural features, such as eyes and noses, are successfully matched between different classes.
@@
 Different matchers.
-Figure 5 presents the results with different feature matchers, showing that the proposed method boosts all of them to match corresponding parts of the cheetah and Siberian husky.
-Figure 5: The results with different feature matchers [3, 23, 17].
-SuperPoint [7] is used for the keypoint detection.
-The introduction of the proposed module increases the matching of eyes and noses between the cheetah and Siberian husky and decreases mismatching in the background.
+Figure 6 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
+Table 1: The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions [20].
+The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image.
+Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC.
+The dense matchers, LoFTR [56] and Efficient LoFTR (ELoFTR) [64], are shown for reference.
+Our method improves the robustness of sparse matchers with a slight reduction in clean accuracy and runtime.
+[…]
+keypoint detector : SuperPoint dense matcher keypoint detector : SuperPoint dense matcher
+LG LG+Ours GS GS+Ours LoFTR ELoFTR LG LG+Ours GS GS+Ours LoFTR ELoFTR
+None (Clean) 80.61 78.42 78.04 75.08 80.93 83.48 80.61 78.42 78.04 75.08 80.93 83.48
+Gaussian Noise 43.09 53.27 45.67 52.24 34.97 35.71 27.54 41.01 41.93 44.98 33.02 36.56
+Shot Noise 43.41 53.82 46.10 51.72 40.00 40.36 32.10 42.35 41.75 43.82 36.53 39.52
+Impulse Noise 44.98 50.11 44.91 50.05 37.73 39.89 35.95 43.67 40.97 42.91 34.92 38.63
+Defocus Blur 32.69 48.35 45.58 49.23 57.22 54.97 18.25 32.13 23.16 30.10 49.40 41.97
+Frosted Glass Blur 33.37 47.88 47.40 49.94 54.48 47.09 34.25 44.80 42.96 46.21 54.53 47.52
+Motion Blur 42.12 53.63 44.76 51.84 55.60 48.12 50.07 53.10 44.27 52.69 52.81 45.10
+Zoom Blur 24.40 31.21 22.87 30.97 23.08 22.54 34.67 40.04 31.18 37.11 31.09 28.49
+Snow 31.51 30.33 25.26 23.21 31.85 37.89 56.24 58.15 47.22 48.53 49.15 54.56
+Frost 32.24 31.71 26.39 24.22 19.23 25.49 62.20 64.87 54.34 57.26 43.17 45.18
+Fog 70.99 73.49 71.19 73.89 67.10 67.74 76.96 78.40 75.18 74.87 74.06 75.36
+Brightness 75.48 75.08 73.03 70.84 75.45 79.83 77.14 76.98 76.78 75.06 78.41 81.26
+Contrast 39.50 38.47 39.47 42.61 53.22 59.78 43.09 45.17 40.44 43.53 39.66 52.76
+Elastic Transform 54.78 66.21 60.37 64.36 58.56 60.50 64.93 68.34 68.02 69.92 68.95 71.49
+Pixelate 67.94 68.24 65.01 64.81 74.77 77.27 66.80 68.43 67.16 65.46 76.21 78.45
+JPEG Compression 29.54 36.97 39.66 44.70 59.05 59.80 47.41 53.82 60.39 60.12 68.14 71.70
+Average 44.40 50.58 46.51 49.64 49.49 50.47 48.51 54.08 50.38 52.84 52.67 53.90
+Time [ms/pair] 46.67 68.43 105.72 128.18 312.54 191.17 45.98 68.37 105.34 127.86 305.01 192.53
 5.2 Robust Image Matching
-[…]
-[…]
-keypoint detector : SuperPoint dense
-LG LG+Ours GS GS+Ours LoFTR
-Gaussian Noise 43.09 53.27 45.67 52.24 34.97
-Shot Noise 43.41 53.82 46.10 51.72 40.00
-Impulse Noise 44.98 50.11 44.91 50.05 37.73
-Defocus Blur 32.69 48.35 45.58 49.23 57.22
-Frosted Glass Blur 33.37 47.88 47.40 49.94 54.48
-Motion Blur 42.12 53.63 44.76 51.84 55.60
-Zoom Blur 24.40 31.21 22.87 30.97 23.08
-Snow 31.51 30.33 25.26 23.21 31.85
-Frost 32.24 31.71 26.39 24.22 19.23
-Fog 70.99 73.49 71.19 73.89 67.10
-Brightness 75.48 75.08 73.03 70.84 75.45
-Contrast 39.50 38.47 39.47 42.61 53.22
-Elastic Transform 54.78 66.21 60.37 64.36 58.56
-Pixelate 67.94 68.24 65.01 64.81 74.77
-Jpeg Compression 29.54 36.97 39.66 44.70 59.05
-Average 44.40 50.58 46.51 49.64 49.49
-Time per pair [msec.] 46.67 68.43 105.72 128.18 312.54
-[…]
-[…]
-keypoint detector : SuperPoint dense
-LG LG+Ours GS GS+Ours LoFTR
-None (Clean) 80.61 78.42 78.04 75.08 80.93
-Gaussian Noise 27.54 41.01 41.93 44.98 33.02
-Shot Noise 32.10 42.35 41.75 43.82 36.53
-Impulse Noise 35.95 43.67 40.97 42.91 34.92
-Defocus Blur 18.25 32.13 23.16 30.10 49.40
-Frosted Glass Blur 34.25 44.80 42.96 46.21 54.53
-Motion Blur 50.07 53.10 44.27 52.69 52.81
-Zoom Blur 34.67 40.04 31.18 37.11 31.09
-Snow 56.24 58.15 47.22 48.53 49.15
-Frost 62.20 64.87 54.34 57.26 43.17
-Fog 76.96 78.40 75.18 74.87 74.06
-Brightness 77.14 76.98 76.78 75.06 78.41
-Contrast 43.09 45.17 40.44 43.53 39.66
-Elastic Transform 64.93 68.34 68.02 69.92 68.95
-Pixelate 66.80 68.43 67.16 65.46 76.21
-Jpeg Compression 47.41 53.82 60.39 60.12 68.14
-Average 48.51 54.08 50.38 52.84 52.67
 We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions.
 To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions.
-We address two cases: (i) both input images are corrupted and (ii) only one of them is corrupted.
-The former shows the robustness against corruptions, and the latter shows the robustness against domain shift (i.e., clean image domain v.s.
+We address two cases: (i) both input images are corrupted, and (ii) only one of them is corrupted.
+The former shows the robustness against corruption, and the latter shows the robustness against domain shift (i.e., clean image domain v.s.
 corrupted image domain).
+Figure 6: The results with different matchers [4, 38, 27].
+SuperPoint [10] is used for the keypoint detection.
+The introduction of our module increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
 Dataset.
 We use image pairs from the MegaDepth-1500 [56] test set as the basis for various experiments.
-MegaDepth-1500 is a set of 1,500 image pairs from the MegaDepth [15] dataset for two popular phototourism destinations.
-MegaDepth consists of 3D reconstructed data from one million images, its camera poses and depth maps are first computed from COLMAP [31] and then refined as ground truth.
+MegaDepth [25] dataset consists of 3D reconstructed data from one million images for two popular phototourism destinations, with camera poses and depth maps computed via COLMAP [50] and refined as ground truth.
 Common Corruptions [20] simulates the 15 types of corruptions that occur when images are captured.
@@
 Metric.
-[…]
+We follow the standard evaluation of image matching based on relative pose estimation (e.g., see [27]).
+From the matching results, the relative pose between two images is computed with RANSAC [13] as follows.
+First, an essential matrix representing the positional transformation between the images is calculated, which is then decomposed into a rotation and a translation.
+The camera pose error is computed by the maximum angular error, and the area under the cumulative error curve (AUC) at $20\tcdegree$ is calculated.
 Robustness against common corruptions.
-Table 1 shows the AUC of matching when input images are both corrupted.
-The results of the dense matching method (i.e., LoFTR) is given for reference.
-As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight increase in the runtime.
-The average AUC even exceeds that of LoFTR.
-The matching methods benefit from the proposed method particularly when images are corrupted by noises and blurs.
+The left side of Tab.
+1 shows the AUC of matching when input images are both corrupted.
+The results of the dense matching method (i.e., LoFTR and Efficient LoFTR) are given for reference.
+As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5% to 10% increase) with a slight increase in the runtime.
+The average AUC even exceeds that of Efficient LoFTR.
+The matching methods benefit from the proposed method, particularly when images are corrupted by noises and blurs.
 We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method.
 These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them.
-See Section B.1 in the supplementary materials for more detailed comparisons.
+See Sec.
+B.1 in the supplementary material for more detailed comparisons.
 Robustness against environmental changes.
 As the objects to be matched are non-identical, they can appear in largely different environments.
-Table 2 shows the AUC when only either of the input images is corrupted and the other is clean.
-As in Tab.
+The right side of Tab.
+1 shows the AUC when only one of the input images is corrupted and the other is clean.
+As in the left side of Tab.
 1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case.
@@
 See Section B.2 in the supplementary material for more detailed comparisons.
-Figure 6: Failure cases. (a) High heatmap scores are given to only a small region of objects to be matched. (b) The structures of objects are clearly different.
 5.3 Limitations
@@
 Further, the large importance scores on the irrelevant object disrupt the matching.
-This may be resolved by using a better visual explanation model [47, 32, 2, 1, 20, 34].
+This may be resolved by using a better visual explanation model [73, 3, 1, 31, 54].
 For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched.
-To address these cases, one may introduce recent vision and language models to complement the visual information with text features.
+To address these cases, one may introduce the vision and language model, such as the segment anything model [24, 44] to complement the visual information with text and part-level mask information.
+Figure 7: Failure cases. (a) High heatmap scores are given to only a small region of objects to be matched. (b) The structures of objects are clearly different.
 6 Conclusion
-In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical as in prior studies.
+In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical, as in prior studies.
 Because of the non-identicality, two objects may be captured in different environments and domains.
 Thus, this task also encompasses robust matching under image corruptions and domain shifts.
-The proposed method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects.
+Our method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects.
 Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
-Acknowledgments
-This work was supported by JSPS KAKENHI Grant Number JP22H03658 and JP22K17962. We thank Satomi Tanaka (Ricoh Company, Ltd.) for her fruitful comments and financial support from Ricoh Company, Ltd..
+Acknowledgement.
+Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP22H03658 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
 Supplementary Material
-A More matching non-identical objects
+Appendix A Additional matching results
 In Sec.
-5.1 of the main text, we evaluated non-identical object matching by inserting the proposed method between SuperPoint [7] and LightGlue [17].
-The matching results for various image pair categories show that the proposed method enhances existing matchers and provides dense and consistent correspondences.
+5.1 of the main text, we evaluated non-identical object matching by inserting our method between SuperPoint [10] and LightGlue [27].
+Here, “Ours” in the results refers to the case where only our weighting module is used, without applying the class similarity adapter.
+The matching results for various image pair show that our method enhances matchers and provides dense and consistent correspondences.
 To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$).
+Additionally, we present more results for complex scenes containing multiple objects in a single image.
 Setup.
 We collected various images with objects from the COCO [26] and ImageNet [9] datasets.
-To prepare drawing or sketch versions of several images, we used the DreamStudio [^4] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [28].
+To prepare drawing or sketch versions of several images, we used the DreamStudio [^3] DreamStudio: https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [47].
 For details of the experiment, refer to Sec.
@@
 A.1 Success cases
-Figure 7 shows the matching results for the same class case.
+Figure A shows the matching results for the same class case.
 As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one.
 The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching.
-Even in the case of class discrepancy, as shown in Fig.
-8, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals.
-Figure 9 shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings).
+Even in the case with class discrepancy, as shown in Fig.
+B, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals.
+Figure C shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings).
+Figure D shows cases where multiple objects are present in an image, demonstrating that our class similarity adapter selects semantically similar object pairs and provides condensed matching between the two objects.
 These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
@@
 There are some cases of failure in non-identical object matching, as shown in Fig.
-10.
-It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller), or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object.
+E.
+It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller) or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object.
 Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
-Figure 7: Matching results in the same class case.
+Figure A: Matching results in the same class case.
 Image pairs show objects of the same class in ImageNet [9] or COCO [26] datasets.
 (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] only finds a small number of correspondence, and many of them are incorrect.
-(b) The proposed method significantly improves the matching, and performs dense and consistent correspondences.
-Figure 8: Matching results in the class discrepancy case.
+(b) The proposed method significantly improves the matching and performs dense and consistent correspondences.
+Figure B: Matching results in the class discrepancy case.
 Image pairs show objects with class discrepancies in ImageNet [9] or COCO [26] datasets.
@@
 (b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
-Figure 9: Matching results in the domain shift case.
+Figure C: Matching results in the domain shift case.
 Image pairs for matching are ImageNet [9] or COCO [26] datasets with common corruptions [20], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [47].
@@
 (b) The proposed method reduces these mismatches and further improves to denser matching between objects.
-Figure 10: Failure cases of non-identical object matching.
+Figure D: Matching results in complex scenes containing multiple objects.
+(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] causes correspondences to scatter across multiple objects.
+(b) The proposed class similarity adapter effectively selects semantically similar object pairs and provides condensed and consistent matching between the relevant parts of the objects.
+Figure E: Failure cases of non-identical object matching.
 (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] fails in matching.
-(b) The proposed method improves several matchings but still leaves many corresponding areas unmatched.
-B Robust image matching details
+(b) The proposed method improves several matching but still leaves many corresponding areas unmatched.
+Appendix B Robust image matching details
+With the same setup in Sec.
+5.2 of the main text, we evaluate the effectiveness of the proposed method for LightGlue [27] and GlueStick [38] on the robustness to image corruptions and domain shift in image matching.
+Here, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$.
+The results of relative pose estimation with common corruptions [20] on one or both of the input images in the MegaDepth-1500 [56] show that the proposed method improves the robustness of LightGlue and GlueStick.
+In all experiments, we used the pretrained sparse matchers LightGlue [27], and GlueStick [38], and the dense matchers LoFTR [56] and EfficientLoFTR [64], downloaded from their official repositories.
+[^4] LightGlue: https://github.com/cvg/LightGlue
+GlueStick: https://github.com/cvg/GlueStick
+LoFTR: https://github.com/zju3dv/LoFTR
+ELoFTR: https://github.com/zju3dv/efficientloftr
+B.1 Robustness against common corruptions
+In this experiment, common corruptions are added to both input images to demonstrate robustness against corruption.
+Table A shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
+As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight decrease in clean accuracy.
+For the Noise category, LightGlue and GlueStick with our method even surpass dense matchers in both cases.
+B.2 Robustness against environmental changes
+In contrast to Sec.
+B.1, in this experiment, common corruptions are added to one of the image pairs to demonstrate robustness against domain shift between images.
+Table B shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
+As the results show, the proposed method is robust to LightGlue and GlueStick, and in particular, significantly improves the AUC of the former for all Noise, Blur, and Digital categories, and the latter for all Noise and Blur categories.
+For the Noise category, LightGlue and GlueStick with our method even exceeds dense matchers in both cases.
+Table A: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions.
+Both images are corrupted with the same type of corruption.
+Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC.
+The results of a dense matchers, LoFTR [56] and Efficient LoFTR (ELoFTR) [64], are shown for reference.
+Our method improves the robustness of sparse matchers with a marginal decrease in clean accuracy.
 […]
-B.1 Robustness against common corruptions.
+keypoint detector : SuperPoint dense matcher
+LG LG+Ours GS GS+Ours LoFTR ELoFTR
+None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18
+Gaussian Noise 27.44 37.28 28.64 35.11 20.07 20.32
+Shot Noise 27.07 39.79 27.81 33.42 24.27 23.89
+Impulse Noise 28.70 37.25 29.10 34.44 22.27 23.75
+Defocus Blur 18.58 24.51 30.12 33.78 43.27 40.33
+Frosted Glass Blur 17.24 27.61 29.16 31.75 39.21 30.43
+Motion Blur 27.08 34.85 27.98 35.46 40.68 38.01
+Zoom Blur 12.50 20.23 12.64 20.72 12.01 13.66
+Snow 18.35 12.56 12.46 10.49 18.94 24.27
+Frost 20.86 18.75 17.64 15.48 11.15 15.67
+Fog 56.32 62.62 55.14 57.24 53.05 59.85
+Brightness 61.10 63.95 58.75 56.51 61.88 67.22
+Contrast 24.33 23.10 24.31 27.45 38.57 44.78
+Elastic Transform 36.76 47.02 41.53 45.52 41.27 42.79
+Pixelate 51.60 60.66 47.60 47.32 62.55 64.85
+JPEG Compression 16.60 27.73 22.35 27.64 43.49 43.50
+Average 29.64 35.86 31.02 34.16 35.51 36.89
 […]
-B.2 Robustness against environmental changes.
+None (Clean) 50.43 49.71 45.92 43.11 52.52 56.38
+Gaussian Noise 14.70 24.44 14.63 21.53 9.85 9.57
+Shot Noise 13.71 23.44 13.44 18.06 12.21 11.78
+Impulse Noise 15.65 21.25 16.90 20.64 10.92 11.77
+Defocus Blur 8.22 11.69 16.59 20.64 27.77 25.67
+Frosted Glass Blur 7.17 18.09 14.58 16.12 24.59 16.17
+Motion Blur 14.72 20.72 15.26 22.34 25.37 27.82
+Zoom Blur 5.38 8.82 6.02 14.16 5.17 8.42
+Snow 9.48 8.32 5.92 3.87 9.91 13.79
+Frost 11.92 11.29 9.92 7.72 6.09 8.75
+Fog 39.01 43.43 35.88 37.58 36.94 48.00
+Brightness 43.60 47.98 41.86 39.60 45.13 50.73
+Contrast 11.69 13.93 12.75 15.29 23.13 29.26
+Elastic Transform 20.43 29.64 24.29 27.28 24.40 25.35
+Pixelate 32.97 41.82 29.86 29.66 46.40 48.41
+JPEG Compression 8.00 15.94 9.71 14.79 27.77 27.45
+Average 17.11 22.72 17.84 20.62 22.38 24.20
+Table B: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions.
+Only one of the input images is corrupted and the other is a clean image.
+Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC.
+The results of dense matchers, LoFTR [56], and Efficient LoFTR (ELoFTR) [64] are shown for reference.
+Our method improves the robustness of sparse matchers with a marginal decrease in clean accuracy.
 […]
+keypoint detector : SuperPoint dense matcher
+LG LG+Ours GS GS+Ours LoFTR ELoFTR
+None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18
+Gaussian Noise 16.63 28.34 28.57 32.42 24.17 27.01
+Shot Noise 19.97 32.50 26.23 29.92 26.58 29.11
+Impulse Noise 23.56 34.74 33.12 36.26 25.67 28.67
+Defocus Blur 8.29 22.43 12.03 18.17 36.85 31.36
+Frosted Glass Blur 19.00 32.78 26.39 29.66 40.40 34.59
+Motion Blur 33.29 39.69 27.88 38.30 39.41 37.15
+Zoom Blur 20.38 26.23 16.76 25.67 21.60 21.63
+Snow 39.74 41.07 30.74 33.00 35.91 43.33
+Frost 47.61 51.78 38.80 41.62 32.97 38.00
+Fog 62.39 64.08 61.11 60.80 60.74 65.26
+Brightness 63.40 61.75 62.69 60.04 65.64 69.24
+Contrast 27.81 33.12 25.86 28.95 30.27 40.35
+Elastic Transform 47.62 56.66 52.17 55.08 53.61 56.20
+Pixelate 50.14 56.68 50.33 48.63 63.14 66.22
+JPEG Compression 29.92 40.31 40.93 40.61 53.58 57.23
+Average 33.98 41.48 35.57 38.61 40.70 43.02
 […]
-[…]
-keypoint detector : SuperPoint dense
-LG LG+Ours GS GS+Ours LoFTR
-None (Clean) 67.89 66.98 64.11 61.94 68.88
-Gaussian Noise 27.44 37.28 28.64 35.11 20.07
-Shot Noise 27.07 39.79 27.81 33.42 24.27
-Impulse Noise 28.70 37.25 29.10 34.44 22.27
-Defocus Blur 18.58 24.51 30.12 33.78 43.27
-Frosted Glass Blur 17.24 27.61 29.16 31.75 39.21
-Motion Blur 27.08 34.85 27.98 35.46 40.68
-Zoom Blur 12.50 20.23 12.64 20.72 12.01
-Snow 18.35 12.56 12.46 10.49 18.94
-Frost 20.86 18.75 17.64 15.48 11.15
-Fog 56.32 62.62 55.14 57.24 53.05
-Brightness 61.10 63.95 58.75 56.51 61.88
-Contrast 24.33 23.10 24.31 27.45 38.57
-Elastic Transform 36.76 47.02 41.53 45.52 41.27
-Pixelate 51.60 60.66 47.60 47.32 62.55
-Jpeg Compression 16.60 27.73 22.35 27.64 43.49
-Average 29.64 35.86 31.02 34.16 35.51
-[…]
-Gaussian Noise 14.70 24.44 14.63 21.53 9.85
-Shot Noise 13.71 23.44 13.44 18.06 12.21
-Impulse Noise 15.65 21.25 16.90 20.64 10.92
-Defocus Blur 8.22 11.69 16.59 20.64 27.77
-Frosted Glass Blur 7.17 18.09 14.58 16.12 24.59
-Motion Blur 14.72 20.72 15.26 22.34 25.37
-Zoom Blur 5.38 8.82 6.02 14.16 5.17
-Snow 9.48 8.32 5.92 3.87 9.91
-Frost 11.92 11.29 9.92 7.72 6.09
-Fog 39.01 43.43 35.88 37.58 36.94
-Brightness 43.60 47.98 41.86 39.60 45.13
-Contrast 11.69 13.93 12.75 15.29 23.13
-Elastic Transform 20.43 29.64 24.29 27.28 24.40
-Pixelate 32.97 41.82 29.86 29.66 46.40
-Jpeg Compression 8.00 15.94 9.71 14.79 27.77
-Average 17.11 22.72 17.84 20.62 22.38
-[…]
-[…]
-keypoint detector : SuperPoint dense
-LG LG+Ours GS GS+Ours LoFTR
-None (Clean) 67.89 66.98 64.11 61.94 68.88
-Gaussian Noise 16.63 28.34 28.57 32.42 24.17
-Shot Noise 19.97 32.50 26.23 29.92 26.58
-Impulse Noise 23.56 34.74 33.12 36.26 25.67
-Defocus Blur 8.29 22.43 12.03 18.17 36.85
-Frosted Glass Blur 19.00 32.78 26.39 29.66 40.40
-Motion Blur 33.29 39.69 27.88 38.30 39.41
-Zoom Blur 20.38 26.23 16.76 25.67 21.60
-Snow 39.74 41.07 30.74 33.00 35.91
-Frost 47.61 51.78 38.80 41.62 32.97
-Fog 62.39 64.08 61.11 60.80 60.74
-Brightness 63.40 61.75 62.69 60.04 65.64
-Contrast 27.81 33.12 25.86 28.95 30.27
-Elastic Transform 47.62 56.66 52.17 55.08 53.61
-Pixelate 50.14 56.68 50.33 48.63 63.14
-Jpeg Compression 29.92 40.31 40.93 40.61 53.58
-Average 33.98 41.48 35.57 38.61 40.70
-[…]
-Gaussian Noise 8.42 20.10 16.51 20.52 16.68
-Shot Noise 10.30 20.52 16.14 19.31 18.14
-Impulse Noise 12.88 24.11 21.41 25.35 17.75
-Defocus Blur 3.22 17.92 5.95 15.89 24.73
-Frosted Glass Blur 8.56 22.42 15.54 18.79 26.87
-Motion Blur 18.95 26.22 13.94 24.36 25.92
-Zoom Blur 9.81 19.27 8.95 17.68 15.03
-Snow 22.94 26.90 17.27 20.56 23.94
-Frost 31.58 37.80 24.13 27.95 23.07
-Fog 44.49 45.00 42.65 41.24 44.04
-Brightness 46.14 45.24 43.91 41.63 48.85
-Contrast 14.98 17.15 14.46 17.53 21.09
-Elastic Transform 29.34 39.29 32.27 35.11 36.41
-Pixelate 31.51 38.02 31.82 30.10 46.26
-Jpeg Compression 15.86 27.55 22.24 21.87 37.35
-Average 20.60 28.50 21.81 25.19 28.41
+None (Clean) 50.43 49.71 45.92 43.11 52.52 56.38
+Gaussian Noise 8.42 20.10 16.51 20.52 16.68 18.54
+Shot Noise 10.30 20.52 16.14 19.31 18.14 19.72
+Impulse Noise 12.88 24.11 21.41 25.35 17.75 19.55
+Defocus Blur 3.22 17.92 5.95 15.89 24.73 21.55
+Frosted Glass Blur 8.56 22.42 15.54 18.79 26.87 22.61
+Motion Blur 18.95 26.22 13.94 24.36 25.92 28.77
+Zoom Blur 9.81 19.27 8.95 17.68 15.03 17.16
+Snow 22.94 26.90 17.27 20.56 23.94 31.62
+Frost 31.58 37.80 24.13 27.95 23.07 29.96
+Fog 44.49 45.00 42.65 41.24 44.04 51.15
+Brightness 46.14 45.24 43.91 41.63 48.85 52.85
+Contrast 14.98 17.15 14.46 17.53 21.09 27.79
+Elastic Transform 29.34 39.29 32.27 35.11 36.41 38.80
+Pixelate 31.51 38.02 31.82 30.10 46.26 49.92
+JPEG Compression 15.86 27.55 22.24 21.87 37.35 40.35
+Average 20.60 28.50 21.81 25.19 28.41 31.36

### v4

@@
-Matching Non-Identical Objects
+Matching Semantically Similar Non-Identical Objects
 Abstract
@@
 This study addresses a novel task of matching such non-identical objects at the pixel level.
-We propose a weighting scheme of descriptors that incorporates semantic information from object detectors into existing sparse image matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects.
-The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo–drawing and image corruptions).
+We propose a weighting scheme of descriptors, Semantic Enhancement Weighting (SEW), that incorporates semantic information from object detectors into existing sparse feature matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects.
+The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo vs.
+drawing and image corruptions).
+The code is available at https://github.com/Circ-Leaf/NIOM.
 1 Introduction
@@
 Thanks to this ability, for example, we can find out our missing cats in a gloomy backstreet by remembering them basking in the sun or assembling components to build a shelf based on their perspective drawing.
-In computer vision words, this needs fine-grained matching of feature points between two objects even with class discrepancy and domain shift (e.g., image corruptions).
+In computer vision terms, this needs fine-grained matching of feature points between two objects, even with class discrepancy and domain shift (e.g., image corruptions).
 The abovementioned matching has been partially addressed in the literature.
-In the image matching [49, 56, 27, 17], objects to be matched are assumed identical and captured from different perspectives.
+In the feature matching [69, 79, 37, 24, 39], objects to be matched are assumed identical and captured from different perspectives.
 This assumption also enables supervised training using dense correspondence obtained from 3D reconstruction from image pairs.
-Recently, Transformer-based dense matching methods have shown great performance [56, 45, 74, 23, 11, 72, 5, 67], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [49, 66, 46, 38, 17, 30, 22].
-In contrast, semantic correspondence models [70, 32, 7] focus on matching semantically similar parts between objects (e.g., tires of two bikes).
+Recently, Transformer-based dense matching methods have shown great performance [79, 64, 103, 30, 18, 101, 8, 95, 31], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [69, 92, 65, 54, 24, 43, 29, 68].
+In contrast, semantic correspondence models [99, 98, 45, 34, 62] focus on matching semantically similar parts between objects (e.g., tires of two bikes).
 They are trained on datasets with sparse part-to-part correspondence annotation and thus do not offer fine-grained matching.
-Figure 1: Non-identical object matching by our plug-and-play module incorporated with the sparse image matching pipeline (here, SuperPoint [10] for keypoint detection and LightGlue [27] for matching).
-(a) design differences, (b) class discrepancy, (c) similar objects, and domain shifts such as (d) drawing v.s.
-photo and (e, f) different image corruptions.
-In this study, we generalize image matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig.
+Figure 1: Non-identical object matching between various objects and image styles is achieved by our plug-and-play method with SuperPoint [16] keypoint detector and LightGlue [37] matcher.
+In this study, we generalize feature matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig.
 1).
 For example, the object pair can be street clocks of different designs or four-legged animals of different species.
-Further, the non-identicality naturally requires robustness, e.g., against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s.
+Further, the non-identicality naturally requires robustness against image corruptions and domain shifts (e.g., photo vs.
 drawing) because two objects may be captured in different environments and domains.
 A crucial challenge of achieving non-identical object matching is that we cannot straightforwardly resort to the supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
-We propose a simple plug-and-play weighting module that extends off-the-shelf sparse matching methods to non-identical object matching.
-Given a pair of images, existing sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints.
-Our weighting module is inserted before the feature matching model and weights the descriptors using semantic information.
-Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [51] to obtain weights.
-Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in the non-identical object matching task (see Fig.
+We propose a plug-and-play module, Semantic Enhancement Weighting (SEW), that extends off-the-shelf matchers to non-identical object matching.
+The sparse matching pipeline first detects keypoints in a pair of images and feeds their descriptors into a matcher.
+SEW is inserted before the matcher and weights descriptors using an object detector and Grad-CAM [73].
+Due to task definition, traditional matching relies on low-level features (e.g., color, edges, texture), which are prone to ambiguity and noise (see Fig.
 2).
-The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features.
-Further, to avoid emphasizing irrelevant objects when an image contains more than one object, we propose Class Similarity Adapter, which determines the most relevant object pairs.
-We leverage the class labels from an object detector and determine object pairs based on their semantic similarity in the text embedding space.
+While these features are still vital for fine-grained matching, our high-level feature weighting improves robustness.
+Further, to handle the case with multiple objects in images, we propose Non-visual Object Pairing, which determines the pairs of similar objects to be matched between images and eliminates mismatching between irrelevant objects.
 Our experiments demonstrate that our approach significantly boosts the robustness of sparse matching methods, including LightGlue [37] and GlueStick [54], in two tasks: non-identical object matching and robust image matching.
-In the former task, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance.
-The sparse matchers alone result in scattered mismatching between non-corresponding areas, but combined with our method, fine-grained and consistent matching is achieved.
-In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [56] dataset under common corruptions [20] in terms of the AUC score of standard relative pose estimation.
+For the former, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance, and further provide a quantitative evaluation with a new annotation-free metric, Triangular Matching Consistency (TMC).
+While sparse matchers alone yield scattered mismatches, our method achieves fine-grained and consistent alignment, as reflected in higher TMC scores.
+For the latter, we evaluate the robustness of matching on the MegaDepth-1500 [79] dataset under common corruptions [27] in terms of the AUC score of standard relative pose estimation.
 Our method outperforms the state-of-the-art sparse matcher under various types of image corruptions.
@@
 To summarize, our contributions are as follows:
-  • We tackle non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images, even under image corruptions and/or between different domains, which can be considered a generation of classical image matching and semantic correspodense.
-  • We propose a simple plug-and-play weighting module that enhances the low-level features using semantic information and successfully extends various sparse matching methods to non-identical object matching.
-  • We present a first evaluation of the robustness in image matching under common corruptions.
-By adding corruptions to both input images, the robustness against corruptions itself is measured.
-By adding them to only one of the input images, the robustness against a domain shift is measured.
-For both cases, our method is shown useful.
+  • We tackle non-identical object matching, a novel task of matching similar but not necessarily identical objects in images, even under image corruptions and/or across different domains, which can be considered a generalization of feature matching and semantic correspondence.
+  • We propose Semantic Enhancement Weighting (SEW), which enhances low-level features in a plug-and-play manner using semantic information and successfully extends various matchers to non-identical object matching.
+  • Our experiments show that under challenging conditions such as class discrepancy, image corruptions, and domain shifts, our method yields fine-grained and consistent correspondences, validated by TMC and pose estimation.
 2 Related Work
-Figure 2: Failure cases of non-identical object matching using the state-of-the-art matcher, LightGlue [27], and the semantic correspondence model, GeoAware-SC [70].
-In semantic correspondence models, some points indicating object parts must be designated during inference to find correspondences, and automating this point selection is essential for feature point matching.
-Comparison of feature matching by GeoAware-SC using two methods: one that randomly extracts 50 points and the other that adaptively extracts 50 points based on the GradCAM [51] heatmap of YOLOv7 [62].
-Red boxes indicate mismatches due to background or object-like textures.
-Orange boxes indicate semantically corresponding regions, where most points align at a part level but not at a pixel level.
-For example, in (d), mismatches occur between the tip of the nose of the cheetah and the nostrils of the husky.
-Image matching [49, 56, 27, 17] is a fundamental task in computer vision with various applications [50, 35, 8, 43, 34, 63, 57, 55] that involves finding the fine-grained correspondence between two images.
+Figure 2: State-of-the-art models of sparse matching (LightGlue [37]) and semantic correspondence (GeoAware-SC [98]).
+The latter requires reference points in one image and finding their correspondence in the target image.
+To this end, 50 reference points are sampled randomly or based on the GradCAM [73] heatmap of YOLOv7 [87] (cf.
+Sec.
+D in the supplementary material).
+Red boxes show mismatches.
+Orange boxes show close matches, which are correct at the part level, but not at the pixel level.
+This was typical for Geo-Aware-SC (e.g., matching between the nose tip of the cheetah and the nostrils of the husky in (d)).
+Further, Geo-Aware-SC was about ten times slower than LightGlue.
+(See Tab.
+B in the supplementary material).
+Image matching [69, 79, 37, 24] is a fundamental task in computer vision with various applications [71, 51, 14, 60, 49, 88, 80, 78, 2, 93] that involves finding the fine-grained correspondence between two images.
 Many image matching methods have two stages.
-A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find correspondences between them.
+A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find matches between them.
 Keypoint detectors.
-Classically, keypoint detection was done using hand-crafted local features such as SIFT [28] and ORB [48].
-Recently, learning-based methods have been developed [10, 18, 14, 61, 41, 12].
-Particularly, SuperPoint [10] and DISK [59] are two popular examples.
-The former is based on convolutional neural networks, and the latter is based on reinforcement learning.
-There is also a fully differentiable and lightweight method with a minimal model configuration [14].
-Unsupervised keypoint detection based on a text-to-image diffusion model [18] also exists.
+Classical detectors such as SIFT [40] and ORB [67] rely on hand-crafted features.
+Recent learning-based methods, including SuperPoint [16] and DISK [84], achieve higher robustness, with lightweight [21] and unsupervised variants [25] also proposed.
 Feature matchers.
@@
 The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers.
-The former runs faster, while the latter gives dense and high accuracy.
-The focus of our study lies in sparse matchers.
-A seminal sparse matcher, SuperGlue [49], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [60] with optimal transport [40].
-The state-of-the-art sparse matcher, LightGlue [27], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously.
-GlueStick [38] utilizes line segments for structural features.
-In contrast to sparse matchers, dense matchers [56, 45, 74, 23, 11, 72, 5, 67] peruse matching accuracy rather than speed.
-Dense matchers input an image and infer matching using all pixels as keypoints.
-Therefore, it does not require a keypoint detector, so it is also called a detector-free model.
-LoFTR [56] proposed the first transformer-based method for dense matching, which achieves highly dense matching.
-Recent studies [36, 71, 29] such as Efficient LoFTR [64], enabling fast and accurate matching by token aggregation, ASpanFormer [5], which controls the attention span, and ASTR [67], which focuses on the role of each pixel.
+Sparse matchers (e.g., SuperGlue [69], which combines Transformer attention with optimal transport; LightGlue [37], which introduces a hierarchical matching structure for real-time accuracy; GlueStick [54], which exploits line segments as structural features) use limited keypoints for efficiency.
+In contrast, dense matchers (e.g., LoFTR [79], DKM [17], ROMA [18]) pursue matching accuracy rather than speed by matching all pixels.
 Semantic correspondence models.
-Semantic Correspondence [15, 16, 33, 6, 2] is a task of identifying a semantically similar point in the target image from the given point in the source image.
-Specifically, it enables the correspondence of parts, such as the mouth, nose, or right eye of dogs between the source and target images.
-The training of a model is conducted using a specialized dataset, such as SPair-71k [33] and PF-PASCAL [15], which employ sparsely annotated part-level correspondences as supervisory data.
-Recently, GeoAware-SC [70] has been proposed as a model that leverages pre-trained features from two foundational models, Stable Diffusion [47] and DINOv2 [37], integrating both geometric and semantic information.
-This model significantly improves accuracy on major benchmarks for semantic correspondence, including the SPair-71k [33] and PF-PASCAL [15] datasets.
-In recent years, various models [39, 52, 69, 58, 21] have been studied.
-For example, CATs++ [7], which combines the local feature aggregation capacity of CNNs with the broad contextual comprehension of transformers; a model [32] employing feature maps from the reverse diffusion process in diffusion models; and unsupervised models [19] that learn semantic correspondence without requiring annotated data.
-While feature matching and semantic correspondence are also the tasks of finding correspondences, our task (i.e., fine-grained matching between non-identical objects) is more challenging.
-First, it is hard to be addressed by the supervised learning framework because (i) manual annotation, as in semantic correspondence datasets is laborious for the pixel-level correspondence, and (ii) automatic annotation by 3D reconstruction, as in feature matching, which assumes an identical object viewed from different perspectives, cannot be done.
-Second, because non-identical objects generally appear with different lighting, image corruptions, or even image domains, the matching method should be designed to be robust to such diversity.
-As shown in Fig.
-2, state-of-the-art feature matching methods and semantic correspondence models are susceptible to such variations.
-Several studies addressed robust matching with specific conditions, such as DarkFeat [17] for low-light and SAM [30] for viewpoint and illumination changes.
-We consider broader changes, including fifteen types of common corruptions [20] and several image domains (e.g., photos and sketches).
+Semantic Correspondence [23, 9, 3, 45] is a task of identifying a semantically similar point in the target image from the given point in the source image, with benchmarks such as SPair-71k [48] and PF-PASCAL [22].
+Recent models include GeoAware-SC [98], which integrates geometric cues with pre-trained features, and unsupervised methods without annotations [26].
+It is also worth noting that semantic correspondence models run significantly slower than sparse matching methods because feature extraction depends on large generative models, such as Stable Diffusion [66].
+For example, it runs roughly 10 times slower in Fig.
+2.
+This study addresses non-identical object matching, unifying the fine-grained alignment of feature matching and the high-level matching of semantic correspondence.
+This makes supervised training technically challenging but also offers unique applications (cf.
+Sec.
+3.2).
+In another view, we aim for matching that runs in real-time and is robust to object identity variations, corruptions, and domain shifts.
+This covers broader robustness than prior studies [63, 91, 5], including DarkFeat [24] and SAM [43] for lighting changes.
+See Sec.
+H in the supplementary for detailed related work.
 3 Non-Identical Object Matching
@@
 3.1 Task formulation
-In the following, we present a more formal problem setup based on a standard matching pipeline.
-Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
-Problem 1 (Non-identical object matching)
+We introduce a more formal setup of non-identical object matching.
+To this end, we first present fine-level image matching (i.e., dense pixel-wise matching, as opposed to coarse-level image matching tasks such as semantic correspondence) as a general task and then discuss feature matching and non-identical object matching as special cases.
+Problem 1 (fine-level image matching)
 Let $\mathcal{X}$} be the image domain.
-Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are the position and the descriptor of $i$-th keypoint, respectively.
+Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in[0,1]^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are its position and descriptor of the $i$-th keypoint.
 We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly.
 Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function.
-The non-identical object matching can be formalized as follows.
+The task can be formalized as follows.
 $\min\quad$ ${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
@@
 This condition is implemented by the last two constraints.
-In the image matching task, where an identical object is captured from different angles in two images, the matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
-Our task, non-identical object matching, covers a broader class of matching.
+In the feature matching task, where an identical object is captured from different angles in two images, matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
+Our task, non-identical object matching covers a broader class of matching.
 Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{I}$ can be characterized by several attributes $(o_{I},y_{I},P_{I},D_{I})$ for $I\in\{A,B\}$.
@@
 If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
-The classical image matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$.
+The classical feature matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$.
 In words, it considers identical objects captured from different camera poses in the same domain.
@@
 Below, we highlight two important special cases, both assuming $P_{A}\neq P_{B}$.
-Figure 3: Pipeline of the matching.
-The keypoint detection and feature matching are done by off-the-shelf models.
-The proposed weighting module computes the heatmap scores of objects and weights the descriptors with this semantic information.
+Figure 3: Non-identical object warping. Homography is estimated from fine-grained matching of city clocks with different styles.
 Class discrepancy $(y_{A}\neq y_{B})$.
-Four-legged animals, such as tigers and cats, are objects of different classes but have similar structures.
-They both have two eyes, four legs, and one tail.
-We humans can find a part-by-part matching between them, even with differences in color, texture, and shape.
+Four-legged animals, such as tigers and cats, belong to different classes but have similar structures (two eyes, four legs, and a tail).
+Despite differences in color, texture, and shape, we humans can still find part-by-part matches.
 However, it is not trivial to design matching methods to find fine-grained correspondence while being robust to detail differences.
-If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
 Domain shift $(o_{A}\equiv o_{B},D_{A}\neq D_{B})$.
@@
 This is a robust fine-grained matching of an identical object in different domains.
-Typical domain shifts include drawings v.s.
-photos and clean images v.s.
+Typical domain shifts include photos vs.
+drawings and clean vs.
 corrupted images.
-For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting.
-Compared to photos, drawings have fewer colors, textures, and shadows.
-Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust.
-If robust matching between different domains is realized, for example, one can design a system that guides a worker to assemble parts from the first-person video by comparing the parts with those in a blueprint.
-3.2 Technical challenges
-Generalization from image matching to non-identical object matching introduces ambiguities and perturbations into matching.
-First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework.
-There is no clear measure of a good matching that can be used for computing a loss to minimize.
-Second, domain shifts by common corruptions or image styles significantly change the low-level features.
-To our knowledge, existing matching methods only use low-level features such as colors, edges, and textures.
-As a consequence, these methods are susceptible to the domain shift, as shown in Fig.
-2.
-The first challenge suggests that it would be better to somehow robustify pretrained image matching models instead of training a model from scratch.
-The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.
+For instance, if $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting, the latter lacks colors, textures, and shadows, causing its descriptors to differ significantly.
+Without sufficient robustness, this discrepancy leads to missing correspondences.
+3.2 Distinction from prior tasks and applications
+Non-identical object matching relates to feature matching (i.e., fine-level perspective alignment) and semantic correspondence (i.e., coarse-level part correspondences between similar objects).
+However, achieving fine-grained and semantic alignment of different objects simultaneously is technically non-trivial and has interesting applications.
+Non-identical object matching is challenging because it does not readily fit into supervised training.
+There is no available dataset with fine-grained human annotation of matching between distinct objects, and the construction would be laborious.
+In contrast, feature matching relies on supervised training with ground-truth homography from 3D reconstruction, while semantic correspondence relies on sparse part-level annotations.
+Yet its standard metric, PCK [48, 22], evaluates only coarse part matches and is therefore inadequate for fine-grained alignment.
+The applications of non-identical object matching are unique and cannot be attained by feature matching and semantic correspondence.
+A fine-grained matching with robustness to class discrepancy allows transferring fine-grained landmark annotations.
+For example, computer vision for animals works on datasets targeting a single animal species, such as facial datasets of dogs [46], cats [47], and cattle [12].
+Non-identical object matching allows transferring their annotations to other animals.
+Its robustness to domain shift can also compare the sketch or blueprint of a product to a real one, which helps workers assemble parts and build it up.
+Figure 3 illustrates a simple example, non-identical object warping, that is not covered by prior tasks.
+Figure 4: Pipeline of the matching.
+The keypoint detection and feature matching are done by off-the-shelf models.
+The proposed plug-and-play module, SEW, computes the heatmap scores of objects and weights the descriptors with this semantic information.
 4 Proposed Method
-In this section, we present a simple plug-and-play weighting module that can turn various sparse matchers to work for non-identical object matching.
-Most sparse matchers, such as [49, 27, 4, 38], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher.
+We propose a plug-and-play module, Semantic Enhancement Weighting (SEW), that can be inserted into various sparse matchers, such as SuperGlue [69], LightGlue [37], and more [7, 54].
+Sparse matchers consist of a pair of a keypoint detector and a feature matcher.
+Our module processes the keypoints before feeding them to the feature matcher.
 Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, The keypoint detector outputs collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$.
@@
 Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
-4.1 Weighting module
-The proposed weighting module uses an object detector and visual explanation model.
+As will be shown in Sec.
+5, this simple pipeline works surprisingly well and adapts pretrained sparse matching models without any training, addressing the technical challenge in non-identical object matching discussed above.
+4.1 Semantic Enhancement Weighting
+The proposed weighting module uses an object detector and a visual explanation model.
 This study used YOLOv7 [87] and Grad-CAM [73].
@@
 Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$.
-Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$.
+Then, the scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$.
+[^1] The max operation outperformed other candidates including mean, sigmoid, and softmax operations in our preliminary experiments.
 Finally, the weighted descriptor $\tilde{\bm{d}}_{i}$ is computed by
@@
 We consider that this is because of the softmax operations used in the subsequent feature matching.
-We discuss the consequence of the weighting of descriptors next.
+We next discuss how the weighting of descriptors impacts the matching.
 4.2 Effect of weights in feature matcher
@@
 $\widetilde{\bm{k}}_{i}^{I}$ $=W_{k}\widetilde{\bm{d}}_{i}^{I}=\alpha_{i}W_{k}\bm{d}_{i}^{I}=\alpha_{i}\bm{k}_{i}^{I},$ (7)
-where $I\in\{A,B\}$.
-Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as [^1] To be precise, this score should be considered as the pre-activation score, and the softmax operation will be subsequently applied.
-The terminology of this paper follows the related studies, e.g., [27].
+where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as
 $\widetilde{a}_{ij}^{I}$ $=\widetilde{\bm{q}}_{i}^{I\top}\mathcal{R}\left(\bm{p}_{j}^{I}-\bm{p}_{i}^{I}\right)\widetilde{\bm{k}}_{j}^{I}$ (8)
@@
 The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher.
-Roughly speaking, nothing that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0$ for between those on the background.
+Roughly speaking, noting that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.25$ for between those on the background.
 This is reasonable because we want nice feature representations for object keypoints for nice matching.
@@
 $\widetilde{a}_{ij}^{AB}=\widetilde{\bm{k}}_{i}^{A\top}\widetilde{\bm{k}}_{j}^{B}=(\alpha^{A}_{i}\alpha^{B}_{j})\cdot a_{ij}^{AB}.$ (11)
-As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ is generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
-4.3 Class Similarity Adapter
-Figure 4: Class Similarity Adapter converts class labels from an object detector into embeddings using the CLIP [42] text encoder and compares them with cosine similarity to identify semantically similar object pairs.
+As discussed earlier, $\alpha_{i}^{A}\alpha_{j}^{B}$ is large between keypoints of the objects, encouraging matching between objects.
+4.3 Non-visual Object Pairing
+We next propose Non-visual Object Pairing for the case where images contain multiple objects.
+In such a case, it is essential to determine which object in one image should be matched with another object in the other image, and the weighting module should exploit the heatmaps of them and exclude those of irrelevant objects.
+Otherwise, the local spurious similarity of objects causes scattered and inconsistent matching; see  Fig.
+5 (bottom right), where a nose of a dog and an ear of a cat match.
+A pair of similar objects for matching can be determined using visual or non-visual features.
+For the former, we can use the feature vectors extracted from an intermediate layer of the object detector.
+This can be done simultaneously during the heatmap computation process.
+For the latter, text embedding (e.g., by the CLIP text encoder) of object labels may be a reasonable choice.
+The text embedding vectors can be pre-computed for all the classes if one pursues fast matching.
+The visual and non-visual features are complementary, and the choice should depend on the use case.
+For example, given (lesser panda, raccoon, giant panda), the visually similar pair is (lesser panda, raccoon), while (lesser panda, giant panda) is more conceptually similar.
+Figure 5: Non-visual Object Pairing converts class labels from an object detector into embeddings using the CLIP [59] text encoder and compares them with cosine similarity to identify semantically similar object pairs.
 The GradCAM [73] heatmap of object detector YOLOv7 [87] generated for the chosen object pair is used for our weighting method, enhancing matching with multiple objects.
-Here, we used the keypoint detector SuperPoint [10] and the feature matcher LightGlue [27].
-We propose the Class Similarity Adapter to address scenarios where multiple objects are present in an image.
-Our weighting module uses all heatmaps of objects detected by the object detector for weighting.
-However, this can needlessly emphasize irrelevant objects, leading to scattered matching from one to multiple objects.
-As shown in  Fig.
-4 (bottom right), relying solely on visual information leads to matching based on similarities in texture, color, and shape, where similar black features such as the nose of the dog and the ear of the cat, are incorrectly matched, causing correspondences to scatter across multiple animals.
-To resolve such confusion in the visual space, we again leverage semantic information given by the object detector.
-Particularly, we map the class labels of detected objects to text embedding space and determine the object pair to be matched based on the cosine similarity between embedding vectors.
+We here provide concrete procedures in the Non-visual Object Pairing based on non-visual features.
 Suppose that an image $\bm{x}_{A}$ contains $n_{A}$ objects, and let $Y_{A}=\{y_{i}^{A}\}_{i=1}^{n_{A}}$ be their class labels.
@@
 Particularly, we used the CLIP text encoder [59].
-The Class Similarity Adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
+The adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
 $(y_{i}^{A},y_{j}^{B})=\underset{y^{A}\in Y_{A},~y^{B}\in Y_{B}}{\text{argmax}}\;\frac{\mathcal{T}(y^{A})^{\top}\mathcal{T}(y^{B})}{\|\mathcal{T}(y^{A})\|\|\mathcal{T}(y^{B})\|},$ (12)
-where $\|\,\cdot\,\|$ denotes the $L_{2}$ norm of vector.
-Then, the weighting module only uses the Grad-CAM heatmaps of the $i$-th object in $\bm{x}_{A}$ and the $j$-th object in $\bm{x}_{B}$.
-Based on this result, the weighting module uses the corresponding heatmaps for weighting the descriptors.
-Figure 4 shows this pipeline.
+[…]
 5 Experiments
-We demonstrate the effectiveness of our method in non-identical object matching tasks, presenting qualitative results on various pairs of non-identical (but similar) objects, followed by quantitative evaluation on classical image matching tasks but under common corruptions [20].
+In this section, we demonstrate the effectiveness of the proposed method in non-identical object matching tasks.
+Section 5.1 presents various qualitative comparisons from state-of-the-art sparse feature matching methods.
+Section 5.2 then introduces Triangular Matching Consistency (TMC), a new annotation-free metric that allows us to quantitatively evaluate non-identical object matching itself.
+Section 5.3 presents quantitative evaluations.
+In the latter experiment, as there are no available datasets providing fine-grained matching between non-identical objects, we examine the matching results between identical objects under 15 types of corruptions, which measures the robustness to domain shifts.
+We emphasize that while semantic correspondence is a conceptually similar task to ours, it does not target fine-grained matching, and thus, our experiment setup (i.e., datasets, baseline methods, and metrics) is based on feature matching literature.
+Figure 6: Typical categories of non-identical object matching.
+(a) A combination of SuperPoint [16] and LightGlue [37] only finds a small number of correspondences, and many of them are incorrect.
+(b) Our method SEW significantly improves the matching.
 Setup.
-In the sparse matching framework, the image matching task is performed by a keypoint detector and a feature matcher.
+The sparse matching pipeline involves a keypoint detector and a feature matcher.
 For the former, we adopted pre-trained SuperPoint [16], extracting 2,048 keypoints.
-For the latter, the extracted keypoints are matched using either pretrained SGMNet [4], GlueStick [38], and the state-of-the-art method, LightGlue [27], with and without the proposed weighting module.
+For the latter, the extracted keypoints are matched using either pretrained SGMNet [7], GlueStick [54], and the state-of-the-art method, LightGlue [37], with and without the proposed method.
 We additionally tested a detector-free dense matcher, LoFTR [79] and Efficient LoFTR [89].
 We downloaded the pretrained models from their official repositories (details provided in the supplementary material).
-All experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
+All experiments used NVIDIA GeForce RTX 3090 GPU and 24GB memory.
 5.1 Matching Non-Identical Objects
-In this experiment, we address a matching of non-identical objects.
-It is worth noting that this task can only be qualitatively evaluated because there is no ground truth matching between non-identical objects.
+Here, we perform the matching of non-identical objects and evaluate them qualitatively.
 Datasets.
@@
 To prepare the drawing version of several images, we used the DreamStudio [^2] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [66].
-Figure 5: Typical categories of non-identical object matching.
-(a) A combination of SuperPoint [10] and LightGlue [27] only finds a small number of correspondence, and many of them are incorrect.
-(b) The proposed method significantly improves the matching.
 Robust matching across non-identical objects.
 We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$).
-For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$).
+For all the cases, we have a perspective difference (i.e., $P_{A}\neq P_{B}$).
 Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in Sec.
@@
 Figure 6 shows the matching results with and without the proposed method in the four cases.
-As can be seen in Fig.
-5(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals.
+As shown in Fig.
+6 (a), SuperPoint and LightGlue alone fail to distinguish objects from the background or match corresponding animal structures.
 In contrast, Fig.
-5(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings.
-Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched front wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles.
+6 (b) shows that the proposed method significantly reduces unreasonable matches (e.g., human and racket, animal eye and mouth) while increasing correct part-to-part matchings.
+Particularly, it (i) increased racket matches while reducing mismatches between people, (ii) aligned corresponding animal parts, (iii) matched front wheels despite domain gaps, and (iv) significantly increased matches despite background and angle variations.
 Notably, structural features, such as eyes and noses, are successfully matched between different classes.
@@
 In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
-Different matchers.
-Figure 6 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
+Different matchers. Figure 7 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
+Figure 7: The results with different matchers [7, 54, 37].
+SuperPoint [16] is used for the keypoint detection.
+The introduction of SEW increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
+5.2 Triangular Matching Consistency
+Direct quantitative evaluation of non-identical object matching is difficult, as fine-grained ground-truth correspondences between non-identical objects are inherently ambiguous.
+We therefore introduce Triangular Matching Consistency (TMC), an annotation-free metric.
+TMC measures the agreement between direct matching $(A,C)$ and one-hop matching via $(A,B)$ and $(B,C)$, where $A,B,C$ are semantically related images.
+By selecting $(A,C)$ to be highly similar, their direct matches provide a reliable reference, and a robust our task method should approximate them even through the one-hop path.
+TMC reports error magnitude (RMSE), accuracy (PCK), and coverage (Recall).
+See Sec.
+E and Fig.
+C for the details.
+Dataset.
+We build an image set of 1,500 triplets for TMC.
+We use animal images from the Wild category of the Animal Faces dataset [11].
+This dataset contains diverse animal face images, including dogs, cats, lions, tigers, and more.
+For each image $A_{i}$, we generate a new view $C_{i}$ of the same object by applying a 0.9× pinch-out geometric transformation.
+Then, another species image $A_{j}$ ($j\neq i$) is used as $B$, forming $(A,B,C)=(A_{i},A_{j},C_{i})$.
+Here, $(A,C)$ are identical-instance pairs with viewpoint change, while $B$ provides a semantically similar but different object.
+This design allows us to use $(A,C)$ as a reliable reference while introducing non-identical variation through $B$, which is essential for evaluating our new task.
+Results.
+Table 1 summarizes the evaluation.
+RMSE measures geometric error, and our module reduces it to about half of each baseline.
+It even outperforms dense detector-free methods.
+PCK measures the proportion of matches within a tolerance.
+It improves at all thresholds, with the largest gains at broader ones and clear benefits even under strict settings.
+This shows that our method enhances consistent matching between corresponding object parts.
+Recall measures how well one-hop correspondences cover the $A{\leftrightarrow}C$ baseline, and it also increases with our module.
+Across all three metrics, it indicates that our method enhances the matching consistency of non-identical objects.
+Table 1: Triangular Matching Consistency (TMC) results for the wild category in the Animal Faces dataset [11].
+Sparse matchers (LightGlue (LG), GlueStick (GS)) with SuperPoint; dense matchers (LoFTR, ELoFTR) are shown for reference.
+Metrics include RMSE (lower is better), PCK, and Recall at distance thresholds of 0.01, 0.05, and 0.10 of the image size.
+Our method SEW shows large improvements across all metrics.
+See Sec.
+F for details.
+[…]
+0.01 0.05 0.10 0.01 0.05 0.10
+Sparse Matcher (with SuperPoint) LG 0.14 54.0 68.9 68.2 70.2 76.9 81.4
+LG+SEW 0.07 73.8 87.8 92.6 73.8 87.7 92.9
+GS 0.16 49.0 62.1 62.3 63.8 71.3 73.0
+GS+SEW 0.08 68.8 81.0 86.6 67.3 82.1 86.3
+Dense Matcher (detector-free) LoFTR 0.19 52.7 61.9 66.8 57.3 62.1 65.7
+ELoFTR 0.14 64.5 73.7 76.7 65.9 73.3 76.9
 Table 2: The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions [27].
 The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image.
-Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC.
+Our method SEW largely improves LightGlue (LG) [37] and GlueStick (GS) [54] for most categories and the average AUC.
 The dense matchers, LoFTR [79] and Efficient LoFTR (ELoFTR) [89], are shown for reference.
@@
 keypoint detector : SuperPoint dense matcher keypoint detector : SuperPoint dense matcher
-LG LG+Ours GS GS+Ours LoFTR ELoFTR LG LG+Ours GS GS+Ours LoFTR ELoFTR
+LG LG+SEW GS GS+SEW LoFTR ELoFTR LG LG+SEW GS GS+SEW LoFTR ELoFTR
 None (Clean) 80.61 78.42 78.04 75.08 80.93 83.48 80.61 78.42 78.04 75.08 80.93 83.48
@@
 5.3 Robust Image Matching
-We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions.
-To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions.
+We next quantitatively evaluate the proposed method on a standard image matching task with common corruptions.
+To our knowledge, image matching methods have only been evaluated clean images in literature, except for a few specialized to low-lighting [24] and thermal imaging [83].
 We address two cases: (i) both input images are corrupted, and (ii) only one of them is corrupted.
-The former shows the robustness against corruption, and the latter shows the robustness against domain shift (i.e., clean image domain v.s.
+The former shows robustness to corruptions, while the latter shows it to domain shifts (i.e., clean image domain vs.
 corrupted image domain).
-Figure 6: The results with different matchers [4, 38, 27].
-SuperPoint [10] is used for the keypoint detection.
-The introduction of our module increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
 Dataset.
-We use image pairs from the MegaDepth-1500 [56] test set as the basis for various experiments.
+We used the MegaDepth-1500 [79] test set because it contains many objects, making it suitable for a fair evaluation of the non-identical object matching task.
+This is not the case with other standard datasets [70, 13, 4, 82, 72].
 MegaDepth [35] dataset consists of 3D reconstructed data from one million images for two popular phototourism destinations, with camera poses and depth maps computed via COLMAP [71] and refined as ground truth.
@@
 2 shows the AUC of matching when input images are both corrupted.
-The results of the dense matching method (i.e., LoFTR and Efficient LoFTR) are given for reference.
-As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5% to 10% increase) with a slight increase in the runtime.
+The results of the dense matching methods (i.e., LoFTR and Efficient LoFTR) are given for reference.
+As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5 % to 10 % increase) with a slight increase in runtime.
 The average AUC even exceeds that of Efficient LoFTR.
-The matching methods benefit from the proposed method, particularly when images are corrupted by noises and blurs.
+The matching methods benefit from the proposed method, particularly when images are corrupted by noise and blur.
 We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method.
-These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them.
+These corruptions introduce edgy geometric patterns, and the matcher finds correspondence between them.
 See Sec.
-B.1 in the supplementary material for more detailed comparisons.
+B.2 in the supplementary material for more detailed comparisons.
 Robustness against environmental changes.
@@
 As in the left side of Tab.
-1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case.
+2, the proposed method boosts the LightGlue and GlueStick for most cases and the average case.
 Interestingly, the results are better for almost all weather categories.
-For brightness, it is the only one of all categories that shows a decrease in accuracy, but the drop is slight.
+Only the brightness category shows a decrease in accuracy, but the drop is slight.
 Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions.
 Heatmap weighting is considered to structurally complement the areas hidden by occlusion.
-The proposed method makes existing matchers robust to environmental changes in general, with a slight precision reduction in clean images.
-See Section B.2 in the supplementary material for more detailed comparisons.
-5.3 Limitations
-We here present failure cases.
-Figure 7 shows typical cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences.
-In the former case, the importance scores take large only on particular parts of the objects.
-Further, the large importance scores on the irrelevant object disrupt the matching.
-This may be resolved by using a better visual explanation model [73, 3, 1, 31, 54].
-For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched.
-To address these cases, one may introduce the vision and language model, such as the segment anything model [24, 44] to complement the visual information with text and part-level mask information.
-Figure 7: Failure cases. (a) High heatmap scores are given to only a small region of objects to be matched. (b) The structures of objects are clearly different.
+Overall, the proposed method enhances robustness to environmental changes, with only a slight drop in precision on clean images.
+See Sec.
+B.3 in the supplementary for details.
 6 Conclusion
-In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical, as in prior studies.
-Because of the non-identicality, two objects may be captured in different environments and domains.
-Thus, this task also encompasses robust matching under image corruptions and domain shifts.
-Our method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects.
-Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
-Acknowledgement.
-Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP22H03658 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
+We addressed a novel task of matching non-identical objects, which are only semantically similar, not necessarily identical.
+Due to the non-identicality, two objects may be captured in different environments and domains.
+Thus, robust matching under image corruptions and domain shifts is also encompassed.
+Our method exploits object detection and visual explanation as high-level features, thereby focusing matching on objects.
+The experiments showed the effectiveness of our method in non-identical object matching using TMC, a new annotation-free metric, and relative pose estimation under various image corruptions.
+Acknowledgement
+Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP23K24914 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
 Supplementary Material
@@
 5.1 of the main text, we evaluated non-identical object matching by inserting our method between SuperPoint [16] and LightGlue [37].
-Here, “Ours” in the results refers to the case where only our weighting module is used, without applying the class similarity adapter.
-The matching results for various image pair show that our method enhances matchers and provides dense and consistent correspondences.
+Here, “SEW” in the results refers to the case where only our weighting module is used without applying the Non-visual Object Pairing.
+The matching results for various image pairs show that our method enhances matchers and provides dense and consistent correspondences.
+The fine-grained correspondence can be applied to tasks such as object warping, as illustrated in Fig.
+A.
 To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$).
@@
 A.1 Success cases
-Figure A shows the matching results for the same class case.
+Figure E shows the matching results for the same class case.
 As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one.
@@
 Even in the case with class discrepancy, as shown in Fig.
-B, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals.
-Figure C shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings).
-Figure D shows cases where multiple objects are present in an image, demonstrating that our class similarity adapter selects semantically similar object pairs and provides condensed matching between the two objects.
+F, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals.
+Figure G shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings).
+Figure H shows cases where multiple objects are present in an image, demonstrating that our Non-visual Object Pairing selects semantically similar object pairs and provides condensed matching between the two objects.
 These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
@@
 There are some cases of failure in non-identical object matching, as shown in Fig.
-E.
+I.
 It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller) or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object.
 Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
-Figure A: Matching results in the same class case.
-Image pairs show objects of the same class in ImageNet [9] or COCO [26] datasets.
-(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] only finds a small number of correspondence, and many of them are incorrect.
-(b) The proposed method significantly improves the matching and performs dense and consistent correspondences.
-Figure B: Matching results in the class discrepancy case.
-Image pairs show objects with class discrepancies in ImageNet [9] or COCO [26] datasets.
-(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence.
-(b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
-Figure C: Matching results in the domain shift case.
-Image pairs for matching are ImageNet [9] or COCO [26] datasets with common corruptions [20], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [47].
-(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds correct matching, but also finds mismatches with many non-correspondence areas.
-(b) The proposed method reduces these mismatches and further improves to denser matching between objects.
-Figure D: Matching results in complex scenes containing multiple objects.
-(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] causes correspondences to scatter across multiple objects.
-(b) The proposed class similarity adapter effectively selects semantically similar object pairs and provides condensed and consistent matching between the relevant parts of the objects.
-Figure E: Failure cases of non-identical object matching.
-(a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] fails in matching.
-(b) The proposed method improves several matching but still leaves many corresponding areas unmatched.
+Figure A: Object warping by homography transformation for fine-grained and consistent matching using our method.
+(b) By warping the object in the right image to the object position in the left image and overlaying the two images, (c) it can be confirmed that the object positions are superimposed, showing consistent alignment.
+Table A: Ablation study on the presence or absence of Non-visual Object Pairing in our method (Weighting Module: SEW, Non-visual Object Pairing: NOP).
+The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions [27].
+The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image.
+NOP selects similar objects between images, but the MegaDepth-1500 dataset assumes that the same object appears in both images.
+Therefore, there is little need to select the object to be matched, resulting in minimal accuracy differences with and without NOP.
+[…]
+keypoint detector : SuperPoint keypoint detector : SuperPoint
+LG LG+SEW LG+SEW+NOP LG LG+SEW LG+SEW+NOP
+None (Clean) 80.61 78.42 78.91 80.61 78.42 78.91
+Gaussian Noise 43.09 53.27 52.21 27.54 41.01 40.00
+Shot Noise 43.41 53.82 52.81 32.10 42.35 41.41
+Impulse Noise 44.98 50.11 48.94 35.95 43.67 42.65
+Defocus Blur 32.69 48.35 47.37 18.25 32.13 31.25
+Frosted Glass Blur 33.37 47.88 47.25 34.25 44.80 44.29
+Motion Blur 42.12 53.63 52.95 50.07 53.10 52.46
+Zoom Blur 24.40 31.21 30.39 34.67 40.04 39.36
+Snow 31.51 30.33 30.40 56.24 58.15 58.00
+Frost 32.24 31.71 31.83 62.20 64.87 64.74
+Fog 70.99 73.49 72.41 76.96 78.40 77.37
+Brightness 75.48 75.08 75.21 77.14 76.98 77.06
+Contrast 39.50 38.47 38.77 43.09 45.17 45.11
+Elastic Transform 54.78 66.21 65.69 64.93 68.34 67.92
+Pixelate 67.94 68.24 68.12 66.80 68.43 68.37
+JPEG Compression 29.54 36.97 36.47 47.41 53.82 53.45
+Average 44.40 50.58 50.05 48.51 54.08 53.56
 Appendix B Robust image matching details
@@
 5.3 of the main text, we evaluate the effectiveness of the proposed method for LightGlue [37] and GlueStick [54] on the robustness to image corruptions and domain shift in image matching.
-Here, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$.
+First, we present an ablation study on the presence or absence of the Non-visual Object Pairing.
+Second, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$.
 The results of relative pose estimation with common corruptions [27] on one or both of the input images in the MegaDepth-1500 [79] show that the proposed method improves the robustness of LightGlue and GlueStick.
@@
 ELoFTR: https://github.com/zju3dv/efficientloftr
-B.1 Robustness against common corruptions
+B.1 Abration on Non-visual Object Pairing
+We conducted an ablation study to assess the impact of our Non-visual Object Pairing (NOP).
+Recall that the NOP assigns object pairs based on the highest cosine similarity of text embeddings (CLIP) derived from their class labels, thereby focusing the Grad-CAM-based weighting on semantically matched objects.
+Table B.1 presents the robustness evaluation against common corruptions applied either to both images or only one of the input images, comPairing results with and without the NOP.
+We find that the accuracy does not vary significantly between these settings.
+This implies that, on the MegaDepth-1500 dataset, where image pairs typically depict the same physical object—our Grad-CAM-based weighting already assigns appropriate focus without needing explicit object selection by the NOP.
+Hence, the NOP’s influence is less pronounced on standard image matching datasets with identical objects.
+In contrast, its benefits become more evident for non-identical object matching tasks, where semantically similar but distinct objects require a more careful pairing mechanism.
+B.2 Robustness against common corruptions
 In this experiment, common corruptions are added to both input images to demonstrate robustness against corruption.
-Table A shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
+Table D shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
 As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight decrease in clean accuracy.
 For the Noise category, LightGlue and GlueStick with our method even surpass dense matchers in both cases.
-B.2 Robustness against environmental changes
+B.3 Robustness against environmental changes
 In contrast to Sec.
-B.1, in this experiment, common corruptions are added to one of the image pairs to demonstrate robustness against domain shift between images.
-Table B shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
+B.2, common corruptions are added to one of the image pairs to demonstrate robustness against domain shift between images.
+Table E shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$.
 As the results show, the proposed method is robust to LightGlue and GlueStick, and in particular, significantly improves the AUC of the former for all Noise, Blur, and Digital categories, and the latter for all Noise and Blur categories.
-For the Noise category, LightGlue and GlueStick with our method even exceeds dense matchers in both cases.
-Table A: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions.
+For the Noise category, LightGlue and GlueStick with our method even exceed dense matchers in both cases.
+Appendix C Other semantic models for weighting
+In our main experiments, we adopted YOLOv7 [87] and Grad-CAM [73] as a fast and practical choice for generating semantic heatmaps.
+Here, we examine alternative semantic models and discuss their effectiveness as weighting sources for descriptors.
+DINOv2 as an alternative.
+We tested DINOv2 [53], a vision foundation model, to generate semantic representations.
+First, we computed principal components of the DINOv2 features for each image to visualize the semantic regions.
+Figure B (a) shows the PCA-based visualization, which captures semantic regions to some extent.
+Then, we generated weight maps by normalizing the PCA scores, as shown in Fig.
+B (b).
+These weight maps were used to reweight keypoint descriptors, following the same procedure as our Grad-CAM-based weighting.
+Effect on matching.
+Figure B (c) illustrates the matching results when the DINOv2-based weight maps were applied.
+Compared to the results using YOLOv7 and Grad-CAM (Fig.
+B (d)), the DINOv2-based weighting did not consistently improve the matching performance.
+In some cases, it even degraded the quality due to overly smooth attention over the image or misaligned semantic focus.
+These results indicate that strong semantic models do not always translate to better weighting for sparse feature matching.
+While DINOv2 performs well in recognition or segmentation tasks, its internal features may not align well with fine-grained keypoint matching needs.
+Figure B: Comparison of weighting.
+(a) Semantic visualization using PCA of DINOv2 features.
+(b) Corresponding weight maps used to reweight descriptors.
+(c) Matching results with DINOv2-based weights.
+(d) Matching results with YOLOv7 and Grad-CAM (baseline).
+The DINOv2-based weights do not necessarily improve the performance compared to the original configuration.
+Table B: Comparison with a semantic correspondence method.
+We report AUC@20 of relative pose estimation using GeoAware-SC [98] and our method (SuperPoint + LightGlue + SEW) on image pairs under common corruptions.
+The left columns show results when both images are corrupted, while the right columns show results when only one image is corrupted (domain shift setup).
+Although both methods perform similarly on clean images, our method consistently outperforms GeoAware-SC under corrupted settings while being more than 10 times faster.
+Common Corruptions corrupted pairs clean / corrupted pairs
+GeoAware-SC Ours GeoAware-SC Ours
+None (clean) 55.28 78.42 55.28 78.42
+gaussian_noise 29.17 53.27 33.57 41.01
+shot_noise 33.82 53.82 36.49 42.35
+impulse_noise 31.39 50.11 34.82 43.67
+defocus_blur 42.32 48.35 26.51 32.13
+glass_blur 39.15 47.88 39.83 44.80
+motion_blur 39.97 53.63 39.36 53.10
+zoom_blur 11.37 21.31 20.78 40.04
+snow 18.44 30.33 35.60 58.15
+frost 10.47 31.71 32.52 64.87
+fog 48.02 73.49 51.31 78.40
+brightness 49.40 75.08 53.48 76.98
+contrast 34.55 38.47 29.58 45.17
+elastic_transform 40.33 66.21 42.75 68.34
+pixelate 48.20 68.24 48.92 68.43
+jpeg_compression 33.06 36.97 42.61 53.82
+Average 33.98 50.82 37.88 54.08
+Time [ms/pair] 710.36 68.43 715.24 68.37
+Appendix D Matching algorithm with Geo-Aware-SC
+In this section, we explain how semantic correspondence is leveraged within the Geo-Aware-SC [98] framework to achieve feature matching.
+We introduce and compare two approaches: a Random Approach and a Heatmap-Based Approach.
+The first method samples pixel locations uniformly, while the second method uses a precomputed heatmap to prioritize semantically meaningful regions.
+D.1 Random sampling approach
+This approach uniformly selects a fixed number of points from the valid image region.
+In this random approach, each image is first resized to a fixed square resolution while preserving its aspect ratio.
+Any remaining space is padded to maintain the square shape and a binary mask is generated to indicate the valid (non-padded) region.
+Next, we apply bilinear interpolation to upsample the feature maps so that each pixel in the resized image can be associated with a distinct feature vector.
+From the valid region, we randomly select a fixed number of pixels for the source image and extract their feature descriptors, which are then normalized.
+The target image undergoes the same resizing and upsampling procedure, but instead of random selection, we use the entire valid region.
+We compute the cosine similarity between every descriptor of the source image and all descriptors of the target image.
+For each descriptor in the source image, the location in the target image with the highest similarity is identified as the match.
+Finally, these correspondences are visualized side-by-side.
+D.2 Heatmap-based sampling approach
+This approach leverages a precomputed semantic heatmap to guide the sampling process, ensuring that pixels in semantically important regions are more likely to be selected.
+The heatmap-based method follows a similar resizing and upsampling procedure, but it incorporates a precomputed heatmap to guide the selection of sample points in the source image.
+The heatmap is aligned to the same resolution as the resized image and normalized.
+Each valid pixel in the source image is associated with a heatmap value, which serves as a sampling probability.
+This probabilistic selection ensures that pixels in regions with higher semantic importance are more likely to be chosen.
+After selecting the sample points based on the heatmap, the descriptors are extracted and normalized just as in the random method.
+The target image also has its feature maps upsampled, and the descriptors are computed for its valid region.
+Each descriptor in the source image is then matched to the most similar descriptor in the target image, allowing important regions indicated by the heatmap to be prioritized during matching.
+D.3 Comparison with Semantic Matching Methods
+To further contextualize the effectiveness of our method, we compared it with GeoAware-SC [98], a representative semantic correspondence method.
+Although semantic matching techniques such as GeoAware-SC are typically designed for high-level alignment of semantically similar regions rather than fine-grained keypoint-level matching, they serve as a useful reference for evaluating robustness under challenging conditions.
+The experimental setup followed that of Sec.
+5.3, focusing on relative pose estimation under common corruptions.
+As presented in Tab.
+B, our method achieves significantly higher AUC@20 scores for both corrupted image pairs (50.82% vs.
+33.98%) and clean-corrupted image pairs (54.08% vs.
+37.88%) compared to GeoAware-SC.
+Moreover, the average inference time per image pair is an order of magnitude faster (68 ms vs.
+710 ms), indicating practical advantages in efficiency.
+Figure C: Triangular Matching Consistency (TMC) overview.
+Given two images $A$ and $C$ of the same object with different viewpoints (where reliable direct correspondences can be defined), and another semantically similar but non-identical object $B$, we compare the direct correspondences between $(A,C)$ with the one-hop correspondences $A{\leftrightarrow}B{\leftrightarrow}C$.
+A method with high TMC produces one-hop matches that closely agree with the direct baseline, thereby enabling annotation-free quantitative evaluation of non-identical object matching.
+Appendix E Triangular Matching Consistency
+We propose Triangular Matching Consistency (TMC), a new annotation-free metric for quantitatively evaluating non-identical object matching.
+Defining ground-truth correspondences between non-identical objects is inherently difficult, as fine-grained pixel-level matches are often ambiguous.
+TMC avoids this limitation by measuring the consistency between the following two matchings: Letting $A,B,C$ be images of non-identical objects, we compute
+  • direct matching of $(A,C)$
+  • one-hop matching through $(A,B)$ and $(B,C)$,
+Images $A,C$ are expected to be very similar, and thus their direct correspondences can serve as a reliable reference.
+A non-identical object matching method is expected to approximate them even via the one-hop matching via semantically similar $B$.
+This design enables quantitative evaluation without annotations and naturally captures cross-object coherence.
+An overview of this procedure is shown in Fig.
+C.
+TMC reports three complementary aspects of performance: (i) error magnitude (RMSE), (ii) accuracy as the fraction of correct correspondences (PCK), and (iii) coverage of the baseline correspondences (Recall).
+We next describe the procedure step by step.
+E.1 Direct and one-hop matchings
+Let $\mathcal{A}$} be the matching algorithm to evaluate. Given images $A,C$, this returns a matching,
+${\mathcal{A}}(A,C)=\quantity{(\bm{p}_{i}^{A},\bm{p}_{i}^{C})}_{i=1}^{M_{1}},$ (13)
+where $\bm{p}_{i}^{A},\bm{p}_{i}^{C}\in[0,1]^{2}$ are the matched position vectors (i.e., normalized coordinates) in $A$ and $C$, respectively, and $M_{1}$ is the number of matched pairs.
+Note that each position vector (say, $\bm{p}_{i}^{A})$ in the matching has an associated descriptor (say, $\bm{d}_{i}^{A})$.
+We denote the one-hop matching via $B$ by
+${\mathcal{A}}(A,C;B)=\quantity{(\widehat{\bm{p}}_{i}^{A},\widehat{\bm{p}}_{i}^{C})}_{i=1}^{M_{2}},$ (14)
+where $M_{2}$ is the number of matched pairs, and this is obtained by combining
+${\mathcal{A}}(A,B)$ $=\quantity{(\bm{q}_{i}^{A},\bm{q}_{i}^{B})}_{i=1}^{N_{1}},$ (15)
+${\mathcal{A}}(B,C)$ $=\quantity{(\bm{r}_{i}^{B},\bm{r}_{i}^{C})}_{i=1}^{N_{2}}.$ (16)
+Particularly, the Hungarian algorithm [50] is applied to pair position vectors in $\bm{q}_{i}^{B}$} and $\bm{r}_{i}^{B}$}, through which the one-hop matching is found.
+For example, if the Hungarian algorithm pairs $(\bm{q}_{i}^{B},\bm{r}_{j}^{B})$, then we have $(\bm{q}_{i}^{A},\bm{r}_{j}^{C})=:(\widehat{\bm{p}}_{i}^{A},\widehat{\bm{p}}_{i}^{C})$.
+The implementation details are as follows.
+Outlier removal.
+The outlier removal by RANSAC-based geometric verification is common in the literature to obtain reliable inlier matches [20, 35, 79, 37].
+In TCM, the removal step is inserted before obtaining Eq.
+13.
+Particularly, given the initial matching between A and C, geometric verification is applied via RANSAC to estimate either a homography or a fundamental matrix, depending on the scene geometry, and only inliers are retained.
+However, the outlier removal is not applied to ${\mathcal{A}}(A,B)$ and ${\mathcal{A}}(B,C)$ because matching between non-identical objects can be regarded as “outliers”; removing them could discard semantically meaningful pairs.
+Proximity used in Hungarian algorithm.
+The image B has two sets of position vectors, $\{\bm{q}_{i}^{B}\}_{i=1}^{N_{1}}$ from ${\mathcal{A}}(A,B)$ and $\{\bm{r}_{i}^{B}\}_{i=1}^{N_{2}}$ from ${\mathcal{A}}(B,C)$.
+The Hungarian algorithm determines the alignment of $M_{2}:=\min\{N_{1},N_{2}\}$ vectors between them by minimizing the cost.
+The cost is defined as follows.
+First, the geometric and descriptor discrepancies are
+$\delta^{p}_{ij}$ $=\|\bm{q}_{i}^{B}-\bm{r}_{j}^{B}\|,$ (17)
+$\delta^{d}_{ij}$ $=\|\bm{e}_{i}^{B}-\bm{f}_{j}^{B}\|,$ (18)
+where $\bm{e}_{i}^{B},\bm{f}_{j}^{B}$ are the descriptors associated with $\bm{q}_{i}^{B},\bm{r}_{j}^{B}$, respectively. The cost is then defined by
+$c_{ij}$ $=\lambda_{p}\,\delta^{p}_{ij}+\lambda_{d}\,\delta^{d}_{ij},$ (19)
+where $\lambda_{p},\lambda_{d}\geq 0$ are hyper-parameters. In our experiments, we set $\lambda_{p}=1.0$ and $\lambda_{d}=0.3$.
+E.2 Measuring matching consistency
+Using the direct and one-hop matching, we now measure the matching consistency by three complementary metrics.
+The consistency is measured using the position vectors of the image $C$.
+Recall that we have $\quantity{\bm{p}_{i}^{C}}_{i=1}^{M_{1}}$ from the direct matching and $\quantity{\widehat{\bm{p}}_{i}^{C}}_{i=1}^{M_{2}}$ from the one-hop matching.
+Let
+$\Delta_{ij}=\|\bm{p}^{C}_{i}-\widehat{\bm{p}}^{C}_{i}\|_{2},\quad i=1,\ldots,M_{1},\ j=1,\ldots,M_{2},$ (20)
+and we introduce three metrics of consistency as follows.
+RMSE (Root Mean Squared Error). RMSE measures the magnitude of geometric error between the direct and one-hop matching.
+$\mathrm{RMSE}$ $=\sqrt{\frac{1}{M_{2}}\sum_{j=1}^{M_{2}}\min_{i}\Delta_{ij}^{2}}.$ (21)
+Noting that the position vectors are normalized, its value ranges in $0,\sqrt{2}]$. Smaller values indicate better consistency.
+PCK (Percentage of Correct Keypoints). PCK [22] measures the proportion of accurate correspondences within a tolerance $\tau>0$ in the normalized plane:
+$\mathrm{PCK}(\tau)$ $=\frac{1}{M_{2}}\sum_{j=1}^{M_{2}}\mathbf{1}[\min_{i}\Delta_{ij}\leq\tau],$ (22)
+where $\mathbf{1}[\,\cdot\,]$ is the indicator function.
+High values at small $\tau$ indicate many highly accurate pairs.
+In our experiments, $\tau$ is set to $0.01$, $0.05$, and $0.1$.
+To reduce sensitivity to any single threshold.
+PCK takes values in $[0,1]$, and larger values indicate better consistency.
+Recall. Recall measures the coverage of the direct matching $A{\leftrightarrow}C$ correspondences by the one-hop correspondences $A{\leftrightarrow}B{\leftrightarrow}C$.
+$\mathrm{Recall}(\tau)$ $=\frac{1}{M_{1}}\sum_{i=1}^{M_{1}}\mathbf{1}[\min_{j}\Delta_{ij}\leq\tau],$ (23)
+In our experiments, $\tau$ is set to $0.01$, $0.05$, and $0.1$. Recall takes values in $[0,1]$, and larger values indicate better coverage.
+Appendix F Additional TMC results
+This section provides additional TMC results that complement the wild-category evaluation presented in Sec.
+5.2.
+We report per-class performance for the dog and cat categories of the Animal Faces dataset [11].
+The evaluation protocol follows the same triplet construction and metrics described in the main paper: error magnitude (RMSE), accuracy (PCK), and coverage (Recall) at distance thresholds of 0.01, 0.05, and 0.10 of the image size.
+Table C shows, as in the wild-category results, our weighting consistently improves all matchers across all metrics.
+Sparse matchers combined with our module achieve substantially lower geometric error and higher accuracy than their baselines, and they also outperform dense detector-free methods in most settings.
+These class-wise results further confirm that our method enhances the robustness and consistency of non-identical object matching.
+Table C: Triangular Matching Consistency (TMC) results for the dog and cat categories in the Animal Faces dataset [11].
+Sparse matchers (LightGlue (LG), GlueStick (GS)) with SuperPoint, and dense matchers (LoFTR, Efficient LoFTR) are shown for reference.
+Metrics include RMSE (lower is better), PCK, and Recall at thresholds of 0.01, 0.05, and 0.10 of the image size.
+[…]
+0.01 0.05 0.10 0.01 0.05 0.10
+Dog Sparse (+SuperPoint) LG 0.20 48.5 60.2 66.1 51.3 62.4 67.8
+LG+SEW 0.15 69.4 77.1 82.0 69.4 76.7 81.9
+GS 0.22 45.0 55.8 60.3 47.9 57.2 61.8
+GS+SEW 0.17 61.8 71.3 77.4 60.4 70.1 76.2
+Dense (no detector) LoFTR 0.26 50.4 59.3 63.7 48.7 56.1 60.9
+ELoFTR 0.20 59.6 68.8 72.5 58.4 67.5 71.2
+Cat Sparse (+SuperPoint) LG 0.13 56.8 72.1 78.4 59.7 74.5 80.1
+LG+SEW 0.08 71.2 84.6 90.4 71.2 84.4 90.2
+GS 0.15 52.7 67.8 74.0 54.9 70.3 75.8
+GS+SEW 0.10 66.5 80.3 86.9 65.1 79.1 86.3
+Dense (no detector) LoFTR 0.17 55.3 70.2 76.1 53.8 69.0 74.8
+ELoFTR 0.12 63.9 78.1 83.4 62.7 77.0 82.6
+Appendix G Limitations
+Figure D shows failure cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences between objects.
+In the former case, the importance scores take large only on particular parts of the objects.
+Further, the large importance scores on the irrelevant object disrupt the matching.
+This may be resolved using better visual explanation methods [102, 6, 1, 44, 77, 94] or vision foundation model [53].
+For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched.
+To address these cases, one may introduce the vision and language model, such as the segment anything model [33, 61] to complement the visual information with text and part-level mask information.
+Appendix H Related Work Details
+Keypoint detectors.
+Classically, keypoint detection was done using hand-crafted local features such as SIFT [40] and ORB [67].
+Recently, learning-based methods have been developed [16, 25, 21, 86, 57, 19].
+Particularly, SuperPoint [16] and DISK [84] are two popular examples.
+The former is based on convolutional neural networks, and the latter is based on reinforcement learning.
+There is also a fully differentiable and lightweight method with a minimal model configuration [21].
+Unsupervised keypoint detection based on a text-to-image diffusion model [25] also exists.
+Feature matchers.
+Feature matchers [90, 96] find a fine-grained correspondence between images.
+The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers.
+The former runs faster, while the latter gives dense and high accuracy.
+The focus of our study lies in sparse matchers.
+A seminal sparse matcher, SuperGlue [69], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [85] with optimal transport [56].
+The state-of-the-art sparse matcher, LightGlue [37], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously.
+GlueStick [54] utilizes line segments for structural features.
+In contrast to sparse matchers, dense matchers [79, 8, 64, 103, 30, 18, 101, 95, 42, 38] pursue matching accuracy rather than speed.
+Dense matchers input an image and infer matching using all pixels as keypoints.
+Therefore, it does not require a keypoint detector, so it is also called a detector-free model.
+LoFTR [79] proposed the first transformer-based method for dense matching, which achieves highly dense matching.
+Recent studies [52, 100, 41] such as Efficient LoFTR [89], enabling fast and accurate matching by token aggregation, and ASTR [95], which focuses on the role of each pixel.
+Figure D: Failure cases: (a) high heatmap scores cover only a small region of objects, (b) the objects have clearly different structures.
+Semantic correspondence models.
+Semantic Correspondence [22, 23, 48, 9, 3, 32, 58, 74] is a task of identifying a semantically similar point in the target image from the given point in the source image.
+Specifically, it enables the correspondence of parts, such as the mouth, nose, or right eye of dogs, between the source and target images.
+The training of a model is conducted using a specialized dataset, such as SPair-71k [48] and PF-PASCAL [22], which employ sparsely annotated part-level correspondences as supervisory data.
+Recently, GeoAware-SC [98] has been proposed to integrate geometric and semantic information by leveraging pre-trained features from Stable Diffusion [66] and DINOv2 [53], boosting semantic correspondence accuracy on major benchmarks [48, 22].
+Additionally, various models [55, 75, 97, 81, 28, 10] have been proposed, including one that uses reverse diffusion feature maps [45] and unsupervised methods that learn semantic correspondence without annotated data [26].
+It is also worth noting that semantic correlation models run significantly slower than sparse matching methods because feature extraction depends on large generative models, such as Stable Diffusion [66].
+For example, it runs roughly 10 times slower in Fig.
+2.
+Figure E: Matching results in the same class case.
+Image pairs show objects of the same class in ImageNet [15] or COCO [36] datasets.
+(a) A combination of SuperPoint (SP) [16] and LightGlue (LG) [37] only finds a small number of correspondences, and many of them are incorrect.
+(b) The proposed method SEW significantly improves the matching and performs dense and consistent correspondences.
+Figure F: Matching results in the class discrepancy case.
+Image pairs show objects with class discrepancies in ImageNet [15] or COCO [36] datasets.
+(a) A combination of SuperPoint (SP) [16] and LightGlue (LG) [37] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence.
+(b) The proposed method SEW finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
+Figure G: Matching results in the domain shift case.
+Image pairs for matching are ImageNet [15] or COCO [36] datasets with common corruptions [27], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [66].
+(a) A combination of SuperPoint (SP) [16] and LightGlue (LG) [37] finds correct matching but also finds mismatches with many non-correspondence areas.
+(b) The proposed method SEW reduces these mismatches and further improves to denser matching between objects.
+Figure H: Matching results in complex scenes containing multiple objects.
+(a) A combination of SuperPoint (SP) [16] and LightGlue (LG) [37] causes correspondences to scatter across multiple objects.
+(b) The proposed Non-visual Object Pairing (NOP) effectively selects semantically similar object pairs and provides condensed and consistent matching between the relevant parts of the objects.
+Figure I: Failure cases of non-identical object matching.
+(a) A combination of SuperPoint (SP) [16] and LightGlue (LG) [37] fails in matching.
+(b) The proposed method SEW improves several matching but still leaves many corresponding areas unmatched.
+Table D: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions.
 Both images are corrupted with the same type of corruption.
@@
 keypoint detector : SuperPoint dense matcher
-LG LG+Ours GS GS+Ours LoFTR ELoFTR
+LG LG+SEW GS GS+SEW LoFTR ELoFTR
 None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18
@@
 Average 17.11 22.72 17.84 20.62 22.38 24.20
-Table B: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions.
+Table E: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions.
 Only one of the input images is corrupted and the other is a clean image.
@@
 keypoint detector : SuperPoint dense matcher
-LG LG+Ours GS GS+Ours LoFTR ELoFTR
+LG LG+SEW GS GS+SEW LoFTR ELoFTR
 None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18